### v1

# Jackknife Empirical Likelihood Method for U Statistics Based on Multivariate Samples and its Applications

Abstract
Empirical likelihood (EL) and its extension via the jackknife empirical likelihood (JEL) method provide robust alternatives to parametric approaches, in the contexts with uncertain data distributions. This paper explores the theoretical foundations and practical applications of JEL in the context of multivariate sample-based U-statistics. In this study we develop the JEL method for multivariate U-statistics with three (or more) samples. This study enhance the JEL methods capability to handle complex data structures while preserving the computation efficiency of the empirical likelihood method. To demonstrate the applications of the JEL method, we compute confidence intervals for differences in VUS measurements which have potential applications in classification problems. Monte Carlo simulation studies are conducted to evaluate the efficiency of the JEL, Normal approximation and Kernel based confidence intervals. These studies validate the superior performance of the JEL approach in terms of coverage probability and computational efficiency compared to other two methods. Additionally, a real data application illustrates the practical utility of the approach. The JEL method developed here has potential applications in dealing with complex data structures.
Empirical Likelihood,
Jackknife,
U Statistics,
Kernel,
Volume under ROC surface,
keywords:
[…]
1 Introduction
Empirical likelihood (EL) has emerged as a powerful tool in statistical inference, particularly valuable when traditional parametric assumptions about data distribution are either unknown or do not hold. EL operates on the principle of maximizing a nonparametric likelihood function constrained by observed data to perform hypothesis tests and construct confidence intervals. Owen’s papers ([15], [16]) have been instrumental in the development and popularization of EL. Furthermore, [17] highlights EL’s ability to automatically shape confidence regions, integrate additional constraints, and preserve the range and transformations inherent in the data. This flexibility and robustness have positioned EL as a cornerstone in modern statistical methodology, offering advantages in scenarios ranging from incomplete data to complex multivariate analyses ([19]).
The jackknife empirical likelihood (JEL) method is an extension of Owen’s foundational work that combines the EL method with the jackknife resampling technique. [7] introduced the JEL method as a modification of EL to deal with non-linear statistics. They illustrated this method using one- and two-sample U-statistics, paving the way for its application in diverse settings. [26] considered testing the equality of two multivariate means using the JEL method. [11] proposed a JEL-based non-parametric inferential procedure for non-smooth estimating functions and demonstrated its application. [18] extended the EL method to general estimating equations based on U-statistics and proposed JEL for vector U-statistics. Using this approach, they developed EL ratio tests for moment-based distribution characteristics. [20] developed a consistent JEL-based test for diagonal symmetry based on energy statistics.
Motivated by these studies, we generalize the JEL method to accommodate $q$-dimensional three (or more)-samples U-statistics. This generalization is needed for a versatile statistical tool capable of handling complex data structures and multiple sample comparisons without relying on distributional assumptions. This extension broadens the applicability of EL and JEL beyond traditional univariate and two-sample scenarios. Additionally, we explain the adaptation of JEL to $q$-dimensional $k$-sample based U-statistics, contribute to a deeper understanding of its theoretical foundations and practical implications in multivariate statistical analysis.
To give an application of described JEL method for three sample based U statistics, we applied it to a practical problem by constructing confidence intervals for the difference in volumes under surfaces (VUS) for two dependent variables. The VUS, introduced by [14], expands upon area under the receiver operating characteristic (ROC) by accommodating three-class diagnostic scenarios (a detailed discussion about VUS is provided in Section 2.1). For an application of the JEL method for $k$-sample-based U statistics, we construct confidence intervals for the difference of two dependent hypervolumes under the ROC manifolds (HUM). ROC curves are generalized to ROC manifolds for multi-category classification, and the analogous extension of the area under the ROC manifold is termed the HUM. For more details on HUM, refer to [10] and Chapter 3 of [25] (a detailed discussion about HUM is provided in Section 5.1).
The paper is organized as follows: In Section 2, we generalize JEL to $q$-dimensional, three-sample-based U-statistics. To provide an application of the described method, we consider the problem of constructing confidence intervals for difference of two dependent VUSs. We use the methods: JEL, U-statistics with asymptotic normality, and the kernel bootstrap to obtain the confidence intervals. Section 3 conducts extensive simulation studies to evaluate effectiveness of the confidence intervals for the difference of the VUSs under the Marshall-Olkin bivariate exponential and Bivariate Pareto models. In Section 4, we consider a real data analysis to demonstrate the findings. In Section 5, we further generalize the JEL U-statistics-based approach to $k\,(>3)$ multivariate sample U-statistics and illustrated the method using HUM.
2 JEL for three multivariate sample
Let $\left(\mathbb{X}_{1},\ldots,\mathbb{X}_{n_{1}}\right),\left(\mathbb{Y}_{1},\ldots,\mathbb{Y}_{n_{2}}\right)$ and $\left(\mathbb{Z}_{1},\ldots,\mathbb{Z}_{n_{3}}\right)$ represent three independent random samples from populations with independent $q$-dimensional distribution functions $F$, $G$ and $H$, respectively, where $\mathbb{X}_{i}=(X_{11},X_{12},\dots,X_{1q})^{T},\;i=1,2,...,n_{1}$, $\mathbb{Y}_{j}=(Y_{11},Y_{12},\dots,Y_{1q})^{T},\;$ $j=1,2,...,n_{2},$ and $\mathbb{Z}_{k}=(Z_{11},Z_{12},\dots,Z_{1q})^{T},\;k=1,2,...,n_{3}$, and $q\geq 2$.
Let $\theta=E[h\left(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},\mathbb{Y}_{j_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}},\mathbb{Z}_{k_{1}}\ldots,\mathbb{Z}_{k_{m_{3}}}\right)]\in\Re$ be the parameter of interest. A three-sample U-statistics of degree $\left(m_{1},m_{2},m_{3}\right)$ with a symmetric kernel $h$ is given by
$U_{n_{1},n_{2},n_{3}}=$ $\frac{1}{{n_{1}\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}\choose m_{3}}}\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}\\ 1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{2}\\ 1\leq k_{1}<\ldots<k_{m_{3}}\leq n_{3}\end{subarray}}h\left(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},\mathbb{Y}_{j_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}},\mathbb{Z}_{k_{1}}\ldots,\mathbb{Z}_{k_{m_{3}}}\right),$ (1)
where $\{i_{1},\ldots,i_{m_{1}}\}$, $\{j_{1},\ldots,j_{m_{2}}\}$ and $\{k_{1},\ldots,k_{m_{3}}\}$ indicates $m_{1}$, $m_{2}$ and $m_{3}$ integers chosen without replacement from sets $\{1,\ldots,n_{1}\}$, $\{1,\ldots,n_{2}\}$ and $\{1,\ldots,n_{3}\}$, respectively.
By using the central limit theorem (see [6]), we have
[…]
Consider
(1) the U statistics based on all observations $U_{n_{1},n_{2},n_{3}}$, as defined by (1);
(2) for $i=1,2,\dots,n_{1},$ the U statistics after deleting observation $X_{i}$, is given by
$U_{n_{1}-1,n_{2},n_{3}}^{(-i)}=\frac{1}{{n_{1}-1\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}\choose m_{3}}}\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}-1\\ i_{l}\neq i,\;l=1,2,\dots,m_{1},\\ 1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{2}\\ 1\leq k_{1}<\ldots<k_{m_{3}}\leq n_{3}\end{subarray}}\hskip-14.45377pth\left(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},\mathbb{Y}_{j_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}},\mathbb{Z}_{k_{1}}\ldots,\mathbb{Z}_{k_{m_{3}}}\right);$ (2)
(3) for $j=1,2,\dots,n_{2},$ the U statistics after deleting observation $Y_{j}$, is given by
$U_{n_{1},n_{2}-1,n_{3}}^{(-j)}=\frac{1}{{n_{1}\choose m_{1}}{n_{2}-1\choose m_{2}}{n_{3}\choose m_{3}}}\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}\\ 1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{2}-1\\ j_{l}\neq j,\;l=1,2,\dots,m_{2},\\ 1\leq k_{1}<\ldots<k_{m_{3}}\leq n_{3}\end{subarray}}\hskip-14.45377pth\left(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},\mathbb{Y}_{j_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}},\mathbb{Z}_{k_{1}}\ldots,\mathbb{Z}_{k_{m_{3}}}\right);$ (3)
(4) for $k=1,2,\dots,n_{3},$ the U statistics after deleting observation $Z_{k}$, is given by
$U_{n_{1},n_{2},n_{3}-1}^{(-k)}=\frac{1}{{n_{1}\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}-1\choose m_{3}}}\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}\\ 1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{2}\\ 1\leq k_{1}<\ldots<k_{m_{3}}\leq n_{3}-1\\ k_{l}\neq k,\;l=1,2,\dots,m_{3}\end{subarray}}\hskip-14.45377pth\left(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},\mathbb{Y}_{j_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}},\mathbb{Z}_{k_{1}}\ldots,\mathbb{Z}_{k_{m_{3}}}\right).$ (4)
The jackknife pseudo-values for individual samples are defined as,
$V_{i,0,0}=n_{1}U_{n_{1},n_{2},n_{3}}-\left(n_{1}-1\right)U_{n_{1}-1,n_{2},n_{3}}^{(-i)},\;\;\forall\;\;i=1,2,\dots,n_{1},$ (5)
$V_{0,j,0}=n_{2}U_{n_{1},n_{2},n_{3}}-\left(n_{2}-1\right)U_{n_{1},n_{2}-1,n_{3}}^{(-j)},\;\;\forall\;\;j=1,2,\dots,n_{2}$ (6)
and
$V_{0,0,k}=n_{3}U_{n_{1},n_{2},n_{3}}-\left(n_{3}-1\right)U_{n_{1},n_{2},n_{3}-1}^{(-k)},\;\;\forall\;\;k=1,2,\dots,n_{3}.$ (7)
It is easy to verify that $\bar{V}_{\cdot,0,0}=\bar{V}_{0,\cdot,0}=\bar{V}_{0,0,\cdot}=U_{n_{1},n_{2},n_{3}}$, where
$\bar{V}_{\cdot,0,0}=\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}V_{i,0,0},\;\;\bar{V}_{0,\cdot,0}=\frac{1}{n_{2}}\sum_{j=1}^{n_{2}}V_{0,j,0}\;\text{ and }\;\bar{V}_{0,0,\cdot}=\frac{1}{n_{3}}\sum_{k=1}^{n_{3}}V_{0,0,k}.$
The jackknife pseudo-values based estimator of $\sigma^{2}$, proposed by [2] (also see [25], [4] and [1]), given as
$\hat{\sigma}^{2}=:\frac{1}{n_{1}\left(n_{1}-1\right)}\sum_{i=1}^{n_{1}}\left(V_{i,0,0}-\bar{V}_{\cdot,0,0}\right)^{2}$ $+\frac{1}{n_{2}\left(n_{2}-1\right)}\sum_{j=1}^{n_{2}}\left(V_{0,j,0}-\bar{V}_{0,\cdot,0}\right)^{2}$
$+\frac{1}{n_{3}\left(n_{3}-1\right)}\sum_{k=1}^{n_{3}}\left(V_{0,0,k}-\bar{V}_{0,0,\cdot}\right)^{2},$ (8)
where $V_{i,0,0}$, $V_{0,j,0}$ and $V_{0,0,k}$ are defined in equations (5), (6) and (7), respectively.
Now, we provide the asymptotic properties of the U-statistics $U_{n_{1},n_{2},n_{3}}$ using the estimator $\hat{\sigma}^{2}$. The following notations are required to describe the theorem as stated below.
$\sigma_{100}^{2}$ $=Var\left[E\left(h\left(\mathbb{X}_{1},\dots,\mathbb{X}_{m_{1}},\mathbb{Y}_{1},\dots,\mathbb{Y}_{m_{2}},\mathbb{Z}_{1},\dots,\mathbb{Z}_{m_{3}}\right)\mid\mathbb{X}_{1}\right)\right],$
$\sigma_{010}^{2}$ $=Var\left[E\left(h\left(\mathbb{X}_{1},\dots,\mathbb{X}_{m_{1}},\mathbb{Y}_{1},\dots,\mathbb{Y}_{m_{2}},\mathbb{Z}_{1},\dots,\mathbb{Z}_{m_{3}}\right)\mid\mathbb{Y}_{1}\right)\right]$
and
$\sigma_{001}^{2}$ $=Var\left[E\left(h\left(\mathbb{X}_{1},\dots,\mathbb{X}_{m_{1}},\mathbb{Y}_{1},\dots,\mathbb{Y}_{m_{2}},\mathbb{Z}_{1},\dots,\mathbb{Z}_{m_{3}}\right)\mid\mathbb{Z}_{1}\right)\right].$
Theorem 2.1.
[…]
(b) Suppose that $\sigma_{100}^{2},\sigma_{010}^{2},\sigma_{001}^{2}>0$ holds. Let $S_{n_{1},n_{2},n_{3}}^{2}=\frac{m_{1}^{2}\sigma_{100}^{2}}{n_{1}}+\frac{m_{2}^{2}\sigma_{010}^{2}}{n_{2}}+\frac{m_{3}^{2}\sigma_{001}^{2}}{n_{3}},$ then,
[…]
and $\;\;\hat{\sigma}^{2}-S_{n_{1},n_{2},n_{3}}^{2}=o_{p}\left(\min\{n_{1},n_{2},n_{3}\}^{-1}\right).$
A two sided $(1-\alpha)$ level CI for $\theta$ based on the above theorem is
$\left(U_{n_{1},n_{2},n_{3}}-z_{\alpha/2}\hat{\sigma},U_{n_{1},n_{2},n_{3}}+z_{\alpha/2}\hat{\sigma}\right),$
where $z_{\alpha}$ is the upper $\alpha$-th percentile point of a standard normal distribution.
For proofs of parts (a) and (b) of the above theorem refer to pages 151-153 of [9]. The proof of the second part of (b) can be easily obtained using Chebyshev’s inequality and is thus omitted (see [4]).
The normal approximation method described above is computationally expensive. Additionally, [7] demonstrated that the JEL method, based on one and two sample U-statistics, outperforms the normal approximation method. Therefore, we develop the JEL method for the above problem to make inferences about the parameter $\theta$. To implement the JEL method, we combine all samples and treat them as a single sample. For this purpose, let $n=n_{1}+n_{2}+n_{3}$ and define $\mathbb{W}_{l}$ as
$\mathbb{W}_{l}=\begin{cases}\mathbb{X}_{l},&\text{ if }l=1,\dots,n_{1}\\ \mathbb{Y}_{l-n_{1}},&\text{ if }l=n_{1}+1,\dots,n_{1}+n_{2}\\ \mathbb{Z}_{l-n_{1}-n_{2}},&\text{ if }l=n_{1}+n_{2}+1,\dots,n\end{cases}.$
Therefore, $(\mathbb{W}_{1},\mathbb{W}_{2},\dots,\mathbb{W}_{n})=(\mathbb{X}_{1},\dots,\mathbb{X}_{n_{1}},\mathbb{Y}_{1},\dots,\mathbb{Y}_{n_{2}},\mathbb{Z}_{1},\dots,\mathbb{Z}_{n_{3}})$, where $\mathbb{W}_{i}^{\prime}s$ are independent but need not be identically distributed. Then, the U-statistics based on single sample $(\mathbb{W}_{1},\mathbb{W}_{2},\dots,\mathbb{W}_{n})$ with degree $m=m_{1}+m_{2}+m_{3}$ is given as follows:
$U_{n}=\frac{1}{{n\choose m}}\mathop{\sum}_{\begin{subarray}{c}1\leq l_{1}<\ldots<l_{m}\leq n\end{subarray}}\tilde{h}\left(\mathbb{W}_{l_{1}},\mathbb{W}_{l_{2}},\ldots,\mathbb{W}_{l_{m}}\right),$ (9)
where
$\tilde{h}\left(\mathbb{W}_{l_{1}},\mathbb{W}_{l_{2}},\ldots,\mathbb{W}_{l_{m}}\right)$
=$=\begin{cases}&\text{if}\;\;1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}\\ c^{*}\;h\left(\mathbb{W}_{i_{1}},\dots,\mathbb{W}_{i_{m_{1}}},\mathbb{W}_{j_{1}},\ldots,\mathbb{W}_{j_{m_{2}}},\mathbb{W}_{k_{1}},\dots,\mathbb{W}_{k_{m_{3}}}\right),&n_{1}+1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{1}+n_{2}\\ &n_{1}+n_{2}+1\leq k_{1}<\ldots<k_{m_{3}}\leq n\\ 0,&\text{otherwise}\end{cases}$
with $c^{*}=\frac{{n\choose m}}{{n_{1}\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}\choose m_{3}}}$. Note that $U_{n_{1},n_{2},n_{3}}=U_{n}$.
Consider the U-statistics after deleting observation $\mathbb{W}_{l}$, from sample $(\mathbb{W}_{1},\mathbb{W}_{2},\dots,\mathbb{W}_{n})$, given by
$U_{n-1}^{(-l)}=\frac{1}{{n-1\choose m}}\mathop{\sum}_{\begin{subarray}{c}1\leq l_{1}<\ldots<l_{m}\leq n\\ l_{i}\neq l,\;i=1,2,\dots,m\end{subarray}}\tilde{h}\left(\mathbb{W}_{l_{1}},\mathbb{W}_{l_{2}},\ldots,\mathbb{W}_{l_{m}}\right),\;\;\forall\;l=1,2,\dots,n.$ (10)
Define the jackknife pseudo-values $V_{l}$, for combined single sample $(\mathbb{W}_{1},\mathbb{W}_{2},\dots,\mathbb{W}_{n})$ as
$V_{l}=nU_{n}-(n-1)U_{n-1}^{(-l)},\;\;\forall\;l=1,2,\dots,n.$ (11)
It is easy to verify that $U_{n}=\frac{1}{n}\sum_{i=1}^{n}V_{i},$ where $V_{i}$ ’s are asymptotically independent random variables ([21]).
The expected values of $V_{l}^{\prime}s$ are given below (provided in Appendix A).
$EV_{l}=\begin{cases}\theta\left(\frac{n}{n-m}\right)\left[n-m-\frac{(n-1)(n_{1}-m_{1})}{n_{1}}\right],&\text{ if }\;l=1,2,\ldots,n_{1}\\ ~{}\\ \theta\left(\frac{n}{n-m}\right)\left[n-m-\frac{(n-1)(n_{2}-m_{2})}{n_{2}}\right],&\text{ if }\;l=n_{1}+1,n_{1}+2,\ldots,n_{1}+n_{2}\\ ~{}\\ \theta\left(\frac{n}{n-m}\right)\left[n-m-\frac{(n-1)(n_{3}-m_{3})}{n_{3}}\right],&\text{ if }\;l=n_{1}+n_{2}+1,\ldots,n\end{cases}.$ (12)
It is easy to verify that $\frac{1}{n}\sum_{l=1}^{n}EV_{l}=\theta$ and, in particular, if $m_{1}=m_{2}=m_{3}$ and $n_{1}=n_{2}=n_{3}$, we get $EV_{l}=\theta$, $\forall\;l=1,2,\ldots,n$.
Finally, we employ the EL method to construct the confidence intervals for $\theta$ by maximizing the empirical likelihood function subject to the constraints based on jackknife pseudo values $V_{l}^{\prime}$s. Let $\mathbf{p}=\left(p_{1},\ldots,p_{n}\right)$ be a probability vector assigning probability $p_{i}$ to each $V_{i}$. The JEL evaluated at $\theta$ becomes
$L(\theta)=\max\left\{\prod_{i=1}^{n}p_{i};\sum_{i=1}^{n}p_{i}=1,\sum_{i=1}^{n}p_{i}\left(V_{i}-EV_{i}\right)=0\right\}.$ (13)
We know that $\prod_{i=1}^{n}p_{i}$ subject to $\sum_{i=1}^{n}p_{i}=1$ attain its maximum at $\frac{1}{n^{n}}$ at $p_{i}=\frac{1}{n}$. Hence the jackknife empirical likelihood ratio for $\theta$ is given by
$R(\theta)=\frac{L(\theta)}{n^{-n}}=\max\left\{\prod_{i=1}^{n}\left(np_{i}\right);\sum_{i=1}^{n}p_{i}=1,\sum_{i=1}^{n}p_{i}\left(V_{i}-EV_{i}\right)=0\right\}.$ (14)
Using Lagrange multipliers method, when
$\min_{1\leq i\leq n}\left(V_{i}-EV_{i}\right)<0<\max_{1\leq i\leq n}\left(V_{i}-EV_{i}\right),$
we obtain
$p_{i}=\frac{1}{n}\frac{1}{1+\lambda\left(V_{i}-EV_{i}\right)},\;\;\forall\;\;i=1,2,\dots,n,$ (15)
where $\lambda$ satisfies
$\frac{1}{n}\sum_{i=1}^{n}\frac{V_{i}-EV_{i}}{1+\lambda\left(V_{i}-EV_{i}\right)}=0.$ (16)
Hence we obtain the jackknife empirical log-likelihood ratio as
$\log R(\theta)=-\sum_{i=1}^{n}\log\left(1+\lambda\left(V_{i}-EV_{i}\right)\right).$
We can use $\log R(\theta)$ for constructing confidence intervals or developing test on $\theta$. For this we need to obtain the asymptotic distribution of $logR(\theta)$. We next derive the limiting distribution of the jackknife empirical log-likelihood ratio as an analogous to Wilk’s theorem. The proof of the following theorem is given in the Appendix B.
Theorem 2.2.
Suppose that following conditions holds:
(a) $E\left(h^{2}\left(\mathbb{X}_{1},\ldots,\mathbb{X}_{m_{1}},\mathbb{Y}_{1},\ldots,\mathbb{Y}_{m_{2}},\mathbb{Z}_{1}\ldots,\mathbb{Z}_{m_{3}}\right)\right)<\infty$ and $\sigma_{100}^{2},\sigma_{010}^{2},\sigma_{001}^{2}>0$,
(b) $0<\lim\inf n_{1}/n_{2}\leq\lim\sup n_{1}/n_{2}<\infty$ and $0<\lim\inf n_{2}/n_{3}\leq\lim\sup n_{2}/n_{3}<\infty$.
Then, as $n\rightarrow\infty,-2\log R(\theta)$ converges in distribution to a $\chi^{2}$ random variable with one degree of freedom.
Using Theorem 2.2, we construct a $(1-\alpha)$ level confidence interval for $\theta$, given by
$CI=\{\theta\;|\;-2\log R(\theta)\leq\chi^{2}_{1,1-\alpha}\},$
where $\chi^{2}_{1,1-\alpha}$ is the $(1-\alpha)$-th percentile point of a $\chi^{2}$ distribution with one degree of freedom.
2.1 Application: Volumes under ROC surfaces
This subsection presents an application of the results obtained above to the Volume Under the ROC Surface (VUS), introduced by [14]. Traditionally confined to binary classifications, ROC curves and their associated Area Under the Curve (AUC) metrics provide critical insights into diagnostic accuracy but are limited in their applicability to multi-class scenarios. The VUS extends AUC by accommodating three-class diagnostic scenarios. This evolution addresses the inherent limitations of binary classifications in medical diagnostics, where conditions often manifest along a spectrum rather than as discrete categories [24]. [25], [4], and [1] considered non-parametric inference for VUS using the JEL method.
Consider levels of three biomarkers used to assess the risk of developing cardiovascular disease: let $X$ be the level of Biomarker 1 (e.g., LDL cholesterol), $Y$ be the level of Biomarker 2 (e.g., HDL cholesterol), and $Z$ be the level of Biomarker 3 (e.g., triglycerides). We are interested in finding the probability that a patient has an LDL cholesterol level $X$ that is lower than their HDL cholesterol level $Y$, and their HDL cholesterol level $Y$ that is lower than their triglyceride level $Z$, i.e., $P(X<Y<Z)$. This probability, which is the VUS, can be interpreted as a measure of how often the biomarker levels are ordered in a specific way that might be associated with different risk levels or health conditions ([5]).
Suppose $X_{i}=\left(X_{i1},X_{i2}\right)^{T},i=1,2,\ldots,n_{1},Y_{j}=\left(Y_{j1},Y_{j2}\right)^{T},j=1,2,\ldots,n_{2}$, and $Z_{k}=\left(Z_{k1},Z_{k2}\right)^{T},k=1,2,\ldots,n_{3},$ represent random samples of three populations having independent joint distribution functions $F,G$ and $H$, respectively.
The difference of two correlated VUS’s can be defined as
$\theta=P\left(X_{11}<Y_{11}<Z_{11}\right)-P\left(X_{12}<Y_{12}<Z_{12}\right).$
In the following part of this subsection, we use three different non-parametric methods and construct confidence intervals for $\theta$.
U-Statistics based Confidence Intervals :
We can rewrite $\theta$ as
$\theta$ $=E\left[I\left(X_{11}<Y_{11}<Z_{11}\right)\right]-E\left[I\left(X_{12}<Y_{12}<Z_{12}\right)\right]$
$=E\left[I\left(X_{11}<Y_{11}<Z_{11}\right)-I\left(X_{12}<Y_{12}<Z_{12}\right)\right],$
where $I\left(x<y<z\right)=\begin{cases}1,&\text{if }\;x<y<z\\ 0,&\text{ otherwise}\end{cases}$.
Consider the symmetric kernel $h^{VUS}$ given by
$h^{VUS}\left(\left(X_{i1},X_{i2}\right),\left(Y_{j1},Y_{j2}\right),\left(Z_{k1},Z_{k2}\right)\right)=I\left(X_{i1}<Y_{j1}<Z_{k1}\right)-I\left(X_{i2}<Y_{j2}<Z_{k2}\right).$
An estimator of $\theta$ can be found as U-statistic with kernel $h$ as
$U_{n_{1},n_{2},n_{3}}^{VUS}=\frac{1}{n_{1}n_{2}n_{3}}\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}\sum_{k=1}^{n_{3}}h^{VUS}\left(\left(X_{i1},X_{i2}\right),\left(Y_{j1},Y_{j2}\right),\left(Z_{k1},Z_{k2}\right)\right).$
The U-statistic $U_{n_{1},n_{2},n_{3}}^{VUS}$ is an unbiased and consistent estimator of $\theta$ (Lehmann, (1951)). Also, using Theorem 2.1, a two sided $(1-\alpha)$ level CI for $\theta$ based on the asymptotic distribution of $U_{n_{1},n_{2},n_{3}}^{VUS}$ is
$\left(U_{n_{1},n_{2},n_{3}}^{VUS}-z_{\alpha/2}\hat{\sigma},U_{n_{1},n_{2},n_{3}}^{VUS}+z_{\alpha/2}\hat{\sigma}\right),$
where $\hat{\sigma}$ is the jackknife pseudo-values based consistent estimator of $Var(U_{n_{1},n_{2},n_{3}}^{VUS})$ (see [2]), as defined by (2).
JEL based Confidence Interval :
Now, we use the Theorem 2.2 to construct JEL based confidence intervals for $\theta$. Here with respect to Theorem 2.2, we have $q=2$ and $m_{1}=m_{2}=m_{2}=1$, and on the same lines as of Section 2, we can define $\log R(\theta)$. Therefore, using Theorem 2.2, the JEL based confidence interval for $\theta$ at level $(1-\alpha)$ is given by
$CI=\{\theta\;|\;-2\log R(\theta)\leq\chi^{2}_{1,1-\alpha}\}.$
Kernel Based Confidence Interval :
To estimate the difference of two VUS’s smoothly [8] proposed the kernel-smoothed version of the difference of two VUS’s as follows
$\hat{\theta}=\frac{1}{n_{1}n_{2}n_{3}}\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}\sum_{k=1}^{n_{3}}\left(\Phi\left(\frac{Y_{1j}-X_{1i}}{\sqrt{h_{1}^{2}+h_{2}^{2}}}\right)\Phi\left(\frac{Z_{1k}-Y_{1j}}{\sqrt{h_{2}^{2}+h_{3}^{2}}}\right)-\Phi\left(\frac{Y_{2j}-X_{2i}}{\sqrt{h_{4}^{2}+h_{5}^{2}}}\right)\Phi\left(\frac{Z_{2k}-Y_{2j}}{\sqrt{h_{5}^{2}+h_{6}^{2}}}\right)\right),$
where $\Phi(\cdot)$ represents the standard normal distribution function and the bandwidths are calculated by
$h_{i}=0.9\min\left\{sd\left(x_{i}\right)\right.,\left.\operatorname{iqr}\left(x_{i}\right)/1.34\right\}n_{i}^{-0.2},\;\;\;\;i=1,2,\dots,6,$
where $sd(\cdot)$ and $\operatorname{iqr}(\cdot)$ represent the standard deviation and the interquartile range. [22] introduced the bandwidth $h_{i}$, which regulates the degree of smoothing. This asymptotic bandwidth selection is effective for a broad range of density functions. Such bandwidths have been utilized by many researchers in diagnostic studies.
To construct the kernel based confidence interval for $\theta$, we employ the non-parametric bootstrap method to obtain the confidence interval. In each bootstrap replication, we find a point estimate $\hat{\theta}$ of $\theta$, that allows us to construct the confidence intervals using the percentile method.
3 Simulation
In this section, we employ Monte Carlo simulation studies to evaluate the finite sample performance of CIs for difference of two VUS’s. The three methods used are the JEL, normal approximation with jackknife variance estimator and kernel based bootstrapping. We assess the performance of each method using two distinct criteria:
(a) Coverage Probability (CP): This represents the probability that the true parameter value falls within the constructed confidence interval (CI). Methods with lower coverage error—defined as the discrepancy between the actual coverage probability and the nominal value are preferred.
(b) Average Length (AL) of CIs: Methods that produce CIs with shorter average lengths are desirable, as shorter intervals provide more precise information about the location of the unknown parameter.
We conduct simulation studies with different sample sizes and simulation is repeated 1,000 time. We use R package “empilik” for finding JEL based confidence intervals. For constructing confidence intervals based on kernel based method we use 100 bootstrap replications. Two bivariate probability models are considered for the simulation studies.
Marshall-Olkin bivariate exponential model (MOBVE($\lambda_{1},\lambda_{2},\lambda_{3}$)) : We consider the Marshall-Olkin bivariate exponential model MOBVE($\lambda_{1},\lambda_{2},\lambda_{3}$), introduced by [13]. The joint cumulative distribution function of this bivariate model is given by
$F_{X_{1},X_{2}}(x_{1},x_{2})=1-\exp[-\lambda_{1}x_{1}-\lambda_{2}x_{2}-\lambda_{3}\max\{x_{1},x_{2}\}],\;\;x_{1}>0,\;x_{2}>0,$
where $\lambda_{i}>0,\;i=1,2,3.$ We generated sample sizes $n_{1}$, $n_{2}$, and $n_{3}$ from three independent models MOBVE($\lambda_{x_{1}},\lambda_{x_{2}},\lambda_{x_{3}}$), MOBVE($\lambda_{y_{1}},\lambda_{y_{2}},\lambda_{y_{3}}$), and MOBVE($\lambda_{z_{1}},\lambda_{z_{2}},\lambda_{z_{3}}$), respectively.
In Table 2, we explored various configurations for sample sizes $(n_{1},n_{2},n_{3})$ and the parameters $(\lambda_{x_{1}},\lambda_{x_{2}},\lambda_{x_{3}};\lambda_{y_{1}},\lambda_{y_{2}},\lambda_{y_{3}};\lambda_{z_{1}},\lambda_{z_{2}},\lambda_{z_{3}})$. We employed the JEL, normal approximation, and kernel-based methods to compute the CP and AL of confidence intervals for $\theta$. This comprehensive investigation aims to evaluate the performance and efficacy of these methods under various parameter settings within the MOBVE framework.
Bivariate Pareto Model : In this part, we explore the Farlie-Gumbel-Morgenstern (FGM) copula based on two independent Pareto distributions with C.D.F.s defined as follows:
$F_{X_{i}}(x_{i})=1-\left(\frac{\lambda_{i}}{x_{i}}\right)^{\alpha_{i}},\;x_{i}\geq\lambda_{i}>0,\;\alpha_{i}>0,\;i=1,2.$
For $-1\leq\theta\leq 1$, the FGM copula, or joint C.D.F., is given by
$C_{\theta}(F_{X_{1}}(x_{1}),F_{X_{2}}(x_{2}))$ $=$ $F_{X_{1}}(x_{1})F_{X_{2}}(x_{2})[1+\theta(1-F_{X_{1}}(x_{1}))(1-F_{X_{2}}(x_{2}))],$
$\qquad\qquad\qquad 0\leq F_{X_{1}}(x_{1}),\,F_{X_{2}}(x_{2})\leq 1$
We refer to this bivariate model as the FGM-based Bivariate Pareto model (FGMPM), denoted as FGMPM($\lambda_{1},\lambda_{2},\alpha_{1},\alpha_{2},\theta$). For simulations, we generated sample sizes $n_{1}$, $n_{2}$, and $n_{3}$ from three FGMPM models: $FGMPM(\lambda_{x_{1}},\lambda_{x_{2}},\alpha_{x_{1}},\alpha_{x_{2}},\theta_{x}),\;FGMPM(\lambda_{y_{1}},\lambda_{y_{2}},\alpha_{y_{1}},\alpha_{y_{2}},\theta_{y}),$ and $FGMPM(\lambda_{z_{1}},\lambda_{z_{2}},\alpha_{z_{1}},\alpha_{z_{2}},\theta_{z}),$ respectively. The samples were generated using the “rCopula” function in the “copula” package in R. Results of the simulation studies for this model are presented in Table 3 for various configurations of sample sizes $(n_{1},n_{2},n_{3})$ and parameters $(\lambda_{x_{1}},\lambda_{x_{2}},\alpha_{x_{1}},\alpha_{x_{2}},\theta_{x};\lambda_{y_{1}},\lambda_{y_{2}},\alpha_{y_{1}},\alpha_{y_{2}},\theta_{y};\lambda_{z_{1}},\lambda_{z_{2}},\alpha_{z_{1}},\alpha_{z_{2}},\theta_{z})$.
The simulations were conducted on a laptop with the following hardware configuration: Processor: Intel(R) Core(TM) i7-7500U CPU @ 2.70GHz, 4 cores; Memory: 8 GB DDR4 RAM; Storage: 240 GB SSD WD green; Operating System: Windows 10, 64-bit.
The approximate time taken by the three methods for obtaining results during simulations is presented in Table 1.
Table 1: Simulation time taken by various methods under two models
Model
Samples Size
$(n_{1},n_{2},n_{3})$
JEL Normal Approx Kernel
MOBVE $(10,10,10)$ 1.6 minutes 1.95 minutes 24 minutes
$(100,100,100)$ 40 hours 69 hours 240 hours
bivariate Pareto $(10,10,10)$ 2.63 minutes 4.6 minutes 50 minutes
$(100,100,100)$ 135 hours 246 hours 656 hours
Table 2: Coverage Probability (CP) and Average length of confidence intervals for the difference two VUS’s at 95% confidence level : Under MOBVE models
$(\lambda_{x_{1}},\lambda_{x_{2}},\lambda_{x_{3}};\lambda_{y_{1}},\lambda_{y_{2}},\lambda_{y_{3}};\lambda_{z_{1}},\lambda_{z_{2}},\lambda_{z_{3}})$ $(n_{1},n_{2},n_{3})\;\;\;$ JEL Normal Approx Kernel
CP (%) AL CP (%) AL CP (%) AL
(1,1,1;1,1,1;1,1,1) (10,10,10) 96.8 .501 97.8 0.384 97 0.392
(20,20,20) 96.5 .332 96.8 0.325 95.8 0.341
(30,30,30) 95.7 .229 97 0.257 96.6 0.264
(50,50,50) 95 .18 96.6 0.183 96 0.195
(70,70,70) 94.4 .126 94 0.124 93.5 0.129
(100,100,100) 95.8 .102 94.2 0.103 94 0.122
(1,2,0;1,1,0;2,1,0) (10,10,10) 97 .564 92.8 0.516 91.8 0.504
(20,20,20) 96.4 .35 93 0.361 92 0.341
(30,30,30) 95.8 .248 92.5 0.239 93 0.25
(50,50,50) 95.4 .187 94 0.187 92.5 0.169
(70,70,70) 95.2 .157 94 0.187 92.5 0.158
(100,100,100) 95.2 .131 94.5 0.130 94 0.135
(1/3,1/3,2/3;2/3,2/3,4/3;1,1,2) (10,10,10) 99.4 .263 99.6 0.234 97.5 0.21
(20,20,20) 97 .169 97.5 0.151 98.5 0.161
(30,30,30) 96 .112 97.5 0.121 98.25 0.125
(50,50,50) 95.6 .085 95.75 0.096 97.5 0.089
(70,70,70) 95.6 .067 94.5 0.069 96.5 0.071
(100,100,100) 95.2 .053 94.8 0.058 96 0.062
(3/5,3/5,2/5;6/5,6/5,4/5;9/5,9/5,6/5) (10,10,10) 99.2 .319 98.75 0.269 96.5 0.235
(20,20,20) 95.2 .2 97 0.197 97.25 0.211
(30,30,30) 97 .146 96.5 0.151 97 0.142
(50,50,50) 96.4 .103 96.5 0.098 97 0.106
(70,70,70) 96 .087 96 0.082 96.5 0.09
(100,100,100) 95.6 .066 96 0.069 96 0.065
(1/19,1/19,18/19;2/19,2/19,36/19;3/19,3/19,54/19) (10,10,10) 98 0.09 86 0.07 88 0.079
(20,20,20) 95.2 0.071 98 0.06 98 0.055
(30,30,30) 97 0.048 99 0.052 94 0.049
(50,50,50) 91 0.034 89 0.034 88.5 0.036
(70,70,70) 90.4 0.027 90 0.028 88 0.029
(100,100,100) 93.6 0.023 91 0.022 90 0.025
Table 3: Coverage Probability (CP) and Average length of confidence intervals for the difference two VUS’s at 95% confidence level : Under Bivariate Pareto models
$(\lambda_{x_{1}},\lambda_{x_{2}},\alpha_{x_{1}},\alpha_{x_{2}},\theta_{x};\lambda_{y_{1}},\lambda_{y_{2}},\alpha_{y_{1}},\alpha_{y_{2}},\theta_{y};\lambda_{z_{1}},\lambda_{z_{2}},\alpha_{z_{1}},\alpha_{z_{2}},\theta_{z})$ $(n_{1},n_{2},n_{3})\;\;\;$ JEL Normal Approx Kernel
CP (%) AL CP (%) AL CP (%) AL
(1,1,1,1,0.5;1,1,1,1,0.5;1,1,1,1,-0.5) (10,10,10) 96 .493 96.3 0.451 98 0.485
(20,20,20) 97 .311 97 0.298 96.6 0.322
(30,30,30) 95 .274 94.3 0.262 94.1 0.281
(50,50,50) 96.8 .192 93.8 0.201 93.7 0.182
(70,70,70) 95.2 .156 94.8 0.153 94 0.159
(100,100,100) 95.4 .109 94.8 0.110 94.2 0.101
(1,2,1,2,0.5;2,1,2,1,0.5;2,2,1,1,-0.5) (10,10,10) 96 .557 91 0.549 89 0.514
(20,20,20) 98 .362 99 0.38 91 0.391
(30,30,30) 92 .268 92 0.306 91 0.321
(50,50,50) 98 .193 99 0.235 93 0.242
(70,70,70) 97 .198 94 0.189 92.5 0.191
(100,100,100) 96 .183 94 0.185 94 0.187
(5,1,1,5,0.2;5,1,2,2,0.5;1,5,5,1,0.9) (10,10,10) 97 .460 90 0.465 88 0.45
(20,20,20) 93 .345 92 0.325 91 0.34
(30,30,30) 94 .275 92 0.295 92 0.29
(50,50,50) 96 .174 97 0.17 98 0.183
(70,70,70) 96 .165 93.5 0.166 94 0.169
(100,100,100) 95.7 .151 94 0.150 94 0.153
(0.5,1,0.5,1,-0.5;1,1,0.5,0.5,-0.5;1,1,5,1,-0.2) (10,10,10) 94 .359 92 0.356 91 0.341
(20,20,20) 98 .212 93 0.201 92 0.204
(30,30,30) 97 .169 93 0.157 92 0.17
(50,50,50) 96 .127 93 0.13 92.5 0.129
(70,70,70) 95.8 .098 93.5 0.095 93 0.097
(100,100,100) 95.5 .076 94 0.075 93.5 0.076
(10,1,0.5,10,-0.9;1,1,15,0.2,-0.1;0.2,1,5,0.1,0.9) (10,10,10) 92 .530 90 0.521 89 0.532
(20,20,20) 98 .317 91 0.29 89 0.299
(30,30,30) 97 .262 91 0.267 91 0.276
(50,50,50) 96 .196 92 0.196 93 0.189
(70,70,70) 96 .163 98 0.165 99 0.17
(100,100,100) 95.8 .123 94 0.12 93.5 0.125
The following observations can be made from Tables 1-3:
1. We observe that for various configurations of parameters of the Marshall-Olkin bivariate exponential and bivariate Pareto models, the average lengths (ALs) of the confidence intervals based on JEL, normal approximation, and kernel-based methods decrease as the sample sizes increase.
2. In most cases, as the sample size grows, the coverage probabilities of confidence intervals generated by the three methods tend to align with the nominal level (95%). This is anticipated because larger sample sizes yield more detailed data insights.
3. From Tables 2-3, we can see that CI based on JEL method achieves closer to 95% coverage across various conditions compared to the other two methods. The average interval lengths are generally better for CIs based on the normal approximation and kernel-based methods for small sample sizes. However, the average lengths are almost same for moderate sample sizes.
4. From Table 1, we can observe that the simulation time taken using the normal approximation with the jackknife variance estimator and kernel-based bootstrapping methods is significantly higher than the time taken by the JEL method.
Therefore, based on the simulation studies, we recommend the JEL method as it requires significantly less time to construct confidence intervals compared to the other two methods and also provides better coverage probability.
4 Data Analysis
[…]
$\theta=P(X_{1}<Y_{1}<Z_{1})-P(X_{2}<Y_{2}<Z_{2}),$
[…]
In total, 883 results were included in this analysis, with sample sizes for the three individual classes being 222, 122, and 539, respectively. We used the JEL, Normal approximation, and Kernel-based methods, as described in Section 2.1, to construct confidence intervals for the difference between two VUSs. The JEL method took 3.4 hours, the Normal approximation method took 4.7 hours, and the Kernel-based method (with 1000 bootstraps) took 52 hours to obtain results for this real data analysis using R software on the previously specified laptop.
[…]
Table 4: Confidence Intervals for $\theta$ at different confidence levels using different methods
Method JEL Normal Approx Kernel
Confidence Interval 90% (0.3124,0.3753) (0.3161,0.3723) (0.3083,0.3743)
Confidence Interval 95% (0.3062,0.3813) (0.3107,0.3776) (0.3006,0.3803)
Confidence Interval 99% (0.3007,0.3866) (0.3002,0.3882) (0.2909,0.3904)
5 JEL for $k\;(>3)$ Multivariate Sample U-statistics
Until now, the results developed in this paper have been restricted to three-sample-based U-statistics. However, in practical applications, there are many cases where U-statistics are based on more than three samples ([10]). To provide an alternative non-parametric method for estimation and testing purposes, in this section, we develop the JEL method for more than three multivariate samples based U statistics. Consider the following independent random samples
$\left(\mathbb{X}_{1}^{(1)},\ldots,\mathbb{X}_{n_{1}}^{(1)}\right),\left(\mathbb{X}_{1}^{(2)},\ldots,\mathbb{X}_{n_{2}}^{(2)}\right),\dots,\left(\mathbb{X}_{1}^{(k)},\ldots,\mathbb{X}_{n_{k}}^{(k)}\right)$
from $k$ independent $q$-dimensional distribution functions $F_{1},F_{2},\dots,F_{k}$, respectively, where
$\mathbb{X}_{i}^{(1)}$ $=(X_{11}^{(1)},X_{12}^{(1)},\dots,X_{1q}^{(1)})^{T},\;i=1,2,...,n_{1},$
$\mathbb{X}_{i}^{(2)}$ $=(X_{11}^{(2)},X_{12}^{(2)},\dots,X_{1q}^{(2)})^{T},\;i=1,2,...,n_{2},$
$\vdots$
$\mathbb{X}_{i}^{(k)}$ $=(X_{11}^{(k)},X_{12}^{(k)},\dots,X_{1q}^{(k)})^{T},\;i=1,2,...,n_{k}.$
Let $\theta=E\left[h\left(\mathbb{X}_{i_{1,1}}^{(1)},\ldots,\mathbb{X}_{i_{1,m_{1}}}^{(1)},\mathbb{X}_{i_{2,1}}^{(2)},\ldots,\mathbb{X}_{i_{2,m_{2}}}^{(2)},\dots,\mathbb{X}_{i_{k,1}}^{(k)}\ldots,\mathbb{X}_{i_{k,m_{k}}}^{(k)}\right)\right]\in\Re$ be the parameter of interest. A $k$-sample U-statistics of degree $\left(m_{1},m_{2},\dots,m_{k}\right)$ with a symmetric kernel $h$ is given by
$U_{n_{1},n_{2},\dots,n_{k}}=$ $d^{*}\hskip-7.22743pt\mathop{\sum\sum\cdots\sum}_{\begin{subarray}{c}1\leq i_{1,1}<\ldots<i_{1,m_{1}}\leq n_{1}\\ 1\leq i_{2,1}<\ldots<i_{2,m_{2}}\leq n_{2}\\ \vdots\;\;\;\\ 1\leq i_{k,1}<\ldots<i_{k,m_{k}}\leq n_{k}\end{subarray}}\hskip-7.22743pth\left(\mathbb{X}_{i_{1,1}}^{(1)},\ldots,\mathbb{X}_{i_{1,m_{1}}}^{(1)},\mathbb{X}_{i_{2,1}}^{(2)},\ldots,\mathbb{X}_{i_{2,m_{2}}}^{(2)},\dots,\mathbb{X}_{i_{k,1}}^{(k)}\ldots,\mathbb{X}_{i_{k,m_{k}}}^{(k)}\right),$
where
$d^{*}=\frac{1}{{n_{1}\choose m_{1}}{n_{2}\choose m_{2}}\cdots{n_{k}\choose m_{k}}},$
and $\{i_{1,1},\ldots,i_{1,m_{1}}\}$, $\{i_{2,1},\ldots,i_{2,m_{2}}\},\dots,\{i_{k,1},\ldots,i_{k,m_{k}}\}$ indicates $m_{1},m_{2},\dots,m_{k}$ integers chosen without replacement from the sets $\{1,\ldots,n_{1}\},\{1,\ldots,n_{2}\},\dots,\{1,\ldots,n_{k}\}$, respectively.
Let $n=\sum_{l=1}^{k}n_{l}$ and $m=\sum_{l=1}^{k}m_{l}$. Define $\mathbb{W}_{l},$ as
$\mathbb{W}_{l}=\begin{cases}\mathbb{X}_{l}^{(1)},&\text{ if }l=1,\dots,n_{1}\\ \mathbb{X}_{l-n_{1}}^{(2)},&\text{ if }l=n_{1}+1,\dots,n_{1}+n_{2}\\ \vdots\\ \mathbb{X}_{l-\sum_{i=1}^{k-1}n_{i}}^{(k)},&\text{ if }l=\sum_{i=1}^{k-1}n_{i}+1,\dots,n\end{cases}$
Therefore,
$(\mathbb{W}_{1},\mathbb{W}_{2},\dots,\mathbb{W}_{n})=\left(\mathbb{X}_{1}^{(1)},\ldots,\mathbb{X}_{n_{1}}^{(1)},\mathbb{X}_{1}^{(2)},\ldots,\mathbb{X}_{n_{2}}^{(2)},\dots,\mathbb{X}_{1}^{(k)},\ldots,\mathbb{X}_{n_{k}}^{(k)}\right).$
In same lines as the three sample case (in Section 2), we define the jackknife pseudo-values for combined single sample $(\mathbb{W}_{1},\mathbb{W}_{2},\dots,\mathbb{W}_{n})$, $V_{l},$ as
$V_{l}=nU_{n}-(n-1)U_{n-1}^{(-l)},\;\;\forall\;\;l=1,2,\dots,n,$ (17)
where $U_{n}$ is the U statistics based on single sample $(\mathbb{W}_{1},\mathbb{W}_{2},\dots,\mathbb{W}_{n})$ with degree $m$ and $U_{n-1}^{(-l)}$ is the U statistics based on sample $(\mathbb{W}_{1},\mathbb{W}_{2},\dots,\mathbb{W}_{n})$, after deleting observation $\mathbb{W}_{l}$. Note that $U_{n}=U_{n_{1},n_{2},\dots,n_{k}}$. It is easy to verify that $U_{n}=\frac{1}{n}\sum_{i=1}^{n}V_{i},$ where $V_{i}$ ’s, as defined by (17), are asymptotically independent random variables ([21]).
The values of $EV_{l}$ are as follows
$EV_{l}=\begin{cases}\theta\left(\frac{n}{n-m}\right)\left[n-m-\frac{(n-1)(n_{1}-m_{1})}{n_{1}}\right],&\text{ if }\;l=1,2,\ldots,n_{1}\\ ~{}\\ \theta\left(\frac{n}{n-m}\right)\left[n-m-\frac{(n-1)(n_{2}-m_{2})}{n_{2}}\right],&\text{ if }\;l=n_{1}+1,n_{1}+2,\ldots,n_{1}+n_{2}\\ \qquad\qquad\qquad\cdot&\\ \qquad\qquad\qquad\cdot&\\ \qquad\qquad\qquad\cdot&\\ \theta\left(\frac{n}{n-m}\right)\left[n-m-\frac{(n-1)(n_{k}-m_{k})}{n_{k}}\right],&\text{ if }\;l=\sum_{i=1}^{k-1}n_{i}+1,\ldots,n\end{cases},$
It is easy to verify that $\frac{1}{n}\sum_{l=1}^{n}EV_{l}=\theta$ and, in particular, if $m_{1}=m_{2}=\dots=m_{k}$ and $n_{1}=n_{2}=\dots=n_{k}$, we get $EV_{l}=\theta$, $l=1,2,\ldots,n$.
Let $\mathbf{p}=\left(p_{1},\ldots,p_{n}\right)$ be a probability vector assigning probability $p_{i}$ to each $V_{i}$. We know that $\prod_{i=1}^{n}p_{i}$ subject to $\sum_{i=1}^{n}p_{i}=1$ attain its maximum at $\frac{1}{n^{n}}$ at $p_{i}=\frac{1}{n}$. Hence the jackknife empirical likelihood ratio for $\theta$ is given by
$R(\theta)=\max\left\{\prod_{i=1}^{n}\left(np_{i}\right);\sum_{i=1}^{n}p_{i}=1,\sum_{i=1}^{n}p_{i}\left(V_{i}-EV_{i}\right)=0\right\}.$
Using Lagrange multipliers method, we obtain the jackknife empirical log-likelihood ratio as
$\log R(\theta)=-\sum_{i=1}^{n}\log\left(1+\lambda\left(V_{i}-EV_{i}\right)\right).$ (18)
Here also we can construct a confidence interval based on Wilks’ theorem. To state the result, we define
$\sigma_{i}^{2}$ $=Var\left[E\left(h\left(\mathbb{X}_{1}^{(1)},\ldots,\mathbb{X}_{m_{1}}^{(1)},\mathbb{X}_{1}^{(2)},\ldots,\mathbb{X}_{m_{2}}^{(2)},\dots,\mathbb{X}_{1}^{(k)},\ldots,\mathbb{X}_{m_{k}}^{(k)}\right)\mid\mathbb{X}_{1}^{(i)}\right)\right],\;\;i=1,2,\dots,k.$
We state the theorem which provides the asymptotic distribution of the jackknife empirical log-likelihood ratio. The proof of the theorem is on the same lines as of Theorem 2.2, and hence omitted.
Theorem 5.1.
Under the following conditions:
(a) $E\left(h^{2}\left(\mathbb{X}_{1}^{(1)},\ldots,\mathbb{X}_{m_{1}}^{(1)},\mathbb{X}_{1}^{(2)},\ldots,\mathbb{X}_{m_{2}}^{(2)},\dots,\mathbb{X}_{1}^{(k)},\ldots,\mathbb{X}_{m_{k}}^{(k)}\right)\right)<\infty$ and $\sigma_{i}^{2}>0,\;\forall\;i=1,2,\dots,k$;
(b) $0<\lim\inf\frac{n_{i-1}}{n_{i}}\leq\lim\sup\frac{n_{i-1}}{n_{i}}<\infty,\;\forall\;i=2,3,\dots,k$,
as $n\rightarrow\infty,-2\log R(\theta)$ converges in distribution to a $\chi^{2}$ random variable with one degree of freedom.
Using Theorem 5.1, the JEL based confidence interval for $\theta$ at $(1-\alpha)$ confidence level is given by
$CI=\{\theta\;|\;-2\log R(\theta)\leq\chi^{2}_{1,1-\alpha}\}.$
Also, the normal approximation based a two-sided confidence interval for $\theta$ at $(1-\alpha)$ confidence level is given by
$\left(U_{n_{1},n_{2},\dots,n_{k}}-z_{\alpha/2}\hat{\tau},U_{n_{1},n_{2},\dots,n_{k}}+z_{\alpha/2}\hat{\tau}\right),$
where $\hat{\tau}^{2}$ is estimator of asymptotic variance of $U_{n_{1},n_{2},\dots,n_{k}}$ that can be obtained in a similar lines as of $\hat{\sigma}^{2}$ in (2).
5.1 Application : Hypervolume Under the ROC Manifold
ROC curves are generalized to ROC manifolds for multi-category classification. The analogous extension of AUC is termed the Hypervolume Under the ROC Manifold (HUM). [10] developed a method for estimating HUM in unordered classifications by using a probabilistic interpretation. They also applied HUM as a criterion for model selection in microarray studies. For more details on HUM, refer to Chapter 3 of [25]. In this section, to illustrate the application of the preceding section, we consider constructing confidence intervals for the difference of two dependent HUMs in k-category classification.
Let
$\left(\mathbb{X}_{1}^{(1)},\ldots,\mathbb{X}_{n_{1}}^{(1)}\right),\left(\mathbb{X}_{1}^{(2)},\ldots,\mathbb{X}_{n_{2}}^{(2)}\right),\dots,\left(\mathbb{X}_{1}^{(k)},\ldots,\mathbb{X}_{n_{k}}^{(k)}\right)$
be $k$ random samples having $k$ independent joint distribution functions $F_{1},F_{2},\dots,F_{k}$, respectively, where
$\mathbb{X}_{i}^{(j)}$ $=(X_{11}^{(j)},X_{12}^{(j)})^{T},\;i=1,2,...,n_{j},\;\;j=1,2,\dots,k.$
Difference of two $k$-category classification dependent HUM is defined as
$\theta=P\left(X_{11}^{(1)}<X_{11}^{(2)}<\cdots<X_{11}^{(k)}\right)-P\left(X_{12}^{(1)}<X_{12}^{(2)}<\cdots<X_{12}^{(k)}\right).$
Therefore, we have
$\theta=E\left[h^{HUM}\left(\mathbb{X}_{1}^{(1)},\mathbb{X}_{1}^{(2)},\dots,\mathbb{X}_{1}^{(k)}\right)\right],$
where
$h^{HUM}\left(\mathbb{X}_{1}^{(1)},\mathbb{X}_{1}^{(2)},\dots,\mathbb{X}_{1}^{(k)}\right)=I_{\left(X_{11}^{(1)}<X_{11}^{(2)}<\cdots<X_{11}^{(k)}\right)}-I_{\left(X_{12}^{(1)}<X_{12}^{(2)}<\cdots<X_{12}^{(k)}\right)}.$
In this case, we have $q=2$ and $m_{1}=m_{2}=\cdots=m_{k}=1,$ and therefore, a $k$-sample U-statistic of degree $\left(1,1,\dots,1\right)$ with a symmetric kernel $h^{HUM}$ is given by
[…]
Using $U_{n}^{HUM}$, we can find jackknife pseudo values $V_{l},\;\forall\;l=1,2,\dots,n,$. Then the log JEL ratio $\log R(\theta)$ can be evaluated using the equation (18). In view of Theorem 5.1, the JEL based $(1-\alpha)$ level confidence interval for $\theta$ is given by
$CI=\{\theta\;|\;-2\log R(\theta)\leq\chi^{2}_{1,1-\alpha}\}.$
And the normal approximation based a two-sided confidence interval for $\theta$ at $(1-\alpha)$ confidence level is given by
$\left(U_{n}^{HUM}-z_{\alpha/2}\hat{\tau},U_{n}^{HUM}+z_{\alpha/2}\hat{\tau}\right).$
6 Concluding remarks
In this paper, we have described the JEL method for $k\,(\geq 3)$ multivariate samples U-statistics. This study opens the door for a wider range of applications in multivariate scenarios. By formalizing the JEL approach for $q$-dimensional $k$-sample-based U-statistics, we have established the asymptotic distribution of the JEL ratio and successfully constructed confidence intervals for unknown parameters.
To demonstrate the applications of our findings, we constructed confidence intervals for differences in two volumes under ROC surfaces (VUSs) and differences in two hypervolumes under the ROC manifolds (HUMs). These applications highlight the practical utility of the JEL method in addressing complex statistical problems. To evaluate the performance of the JEL method in comparison with other non-parametric approaches, we conducted extensive simulation studies under various models, including the Marshall-Olkin bivariate exponential and Bivariate Pareto distributions. Our results showed the superior performance of the JEL method in terms of coverage probability and computational complexity when compared to the normal approximation using jackknife pseudo-values based variance estimator and kernel-based methods.
Additionally, we conducted a real data analysis using an Alzheimer’s disease dataset to illustrate the practical benefits of the JEL method in real-world scenarios. Overall, this study highlights the superior performance, versatility, lower computational cost and practical benefits of the JEL method, recommending its use in statistical practices that require reliable nonparametric inference for complex data structures.
Furthermore, the results of this paper facilitate hypothesis testing using the JEL method, based on the established asymptotic distribution of the JEL ratio. Previous studies, including [23], [12], and Chapter 8 of [3], have explored testing problems based on U-statistics. The JEL method we developed in this paper can be effectively applied to these types of problems.
Appendix A: For Obtaining $EV_{l}$
To obtain expression for $EV_{l}$, we firstly simplify $V_{l}^{\prime}s$.
For simplicity, denote
$\bar{h}_{i_{1},\dots,i_{m_{1}},j_{1},\dots,j_{m_{2}},k_{1},\dots,k_{m_{3}}}\hskip-7.22743pt=h(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},\mathbb{Y}_{j_{1}-n_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}-n_{1}},\mathbb{Z}_{k_{1}-n_{1}-n_{2}}\ldots,\mathbb{Z}_{k_{m_{3}}-n_{1}-n_{2}}).$
For $l=1,2,\dots,n$, we have
$V_{l}=$ $nU_{n}-(n-1)U_{n-1}^{(-l)}$
$=$ $nU_{n}-\frac{(n-1)}{{n-1\choose m}}\mathop{\sum}_{\begin{subarray}{c}1\leq l_{1}<\ldots<l_{m}\leq n\\ l_{i}\neq l,\;i=1,2,\dots,m\end{subarray}}\tilde{h}\left(\mathbb{W}_{l_{1}},\mathbb{W}_{l_{2}},\ldots,\mathbb{W}_{l_{m}}\right)$
$=$ $nU_{n}-\frac{(n-1)}{{n-1\choose m}}\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\\ <j_{1}<\ldots<j_{m_{2}}\\ <k_{1}<\ldots<k_{m_{3}}\leq n-1\\ i_{i}\leq n_{1}<j_{j}\leq n_{1}+n_{2}<k_{k},\\ \forall\;i=1,\dots,m_{1},\;j=1,\dots,m_{2},\;k=1,\dots,m_{3}\\ i_{i}=j_{i}=k_{i}\neq l,\;i=1,2,\dots,m\end{subarray}}\frac{{n\choose m}}{{n_{1}\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}\choose m_{3}}}\bar{h}_{i_{1},\dots,i_{m_{1}},j_{1},\dots,j_{m_{2}},k_{1},\dots,k_{m_{3}}}$
$=$ $nU_{n}-\frac{n(n-1)}{n-m}\frac{1}{{n_{1}\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}\choose m_{3}}}\Bigg{[}\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}-1\\ i_{i}\neq l,\;i=1,2,\dots,m_{1}\\ n_{1}+1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{1}+n_{2}\\ n_{1}+n_{2}+1\leq k_{1}<\ldots<k_{m_{3}}\leq n\end{subarray}}\hskip-14.45377ptI_{[1,n_{1}]}(l)\,\bar{h}_{i_{1},\dots,i_{m_{1}},j_{1},\dots,j_{m_{2}},k_{1},\dots,k_{m_{3}}}$
$\qquad\qquad\qquad+\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}\\ n_{1}+1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{1}+n_{2}-1\\ j_{j}\neq l,\;j=1,2,\dots,m_{2}\\ n_{1}+n_{2}+1\leq k_{1}<\ldots<k_{m_{3}}\leq n\end{subarray}}I_{[n_{1}+1,n_{1}+n_{2}]}(l)\,\bar{h}_{i_{1},\dots,i_{m_{1}},j_{1},\dots,j_{m_{2}},k_{1},\dots,k_{m_{3}}}$
$\qquad\qquad\qquad+\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}\\ n_{1}+1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{1}+n_{2}\\ n_{1}+n_{2}+1\leq k_{1}<\ldots<k_{m_{3}}\leq n-1\\ k_{k}\neq l,\;k=1,2,\dots,m_{3}\end{subarray}}I_{[n_{1}+n_{2}+1,n]}(l)\,\bar{h}_{i_{1},\dots,i_{m_{1}},j_{1},\dots,j_{m_{2}},k_{1},\dots,k_{m_{3}}}\Bigg{]},$ (19)
where $I_{[a,b]}(x)$ is a indicator function, defined as $I_{[a,b]}(x)=\begin{cases}1,&\text{if }a\leq x\leq b\\ 0,&\text{otherwise}\end{cases}.$
Since $U_{n_{1},n_{2},n_{3}}=U_{n}$ and combining (2) and (5), (3) and (6), and (4) and (7), we have the following expressions,
(1) for $i=1,2,\dots,n_{1},$
$\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}-1\\ i_{l}\neq i,\;l=1,2,\dots,m_{1}\\ 1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{2}\\ 1\leq k_{1}<\ldots<k_{m_{3}}\leq n_{3}\end{subarray}}h(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},$ $\mathbb{Y}_{j_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}},\mathbb{Z}_{k_{1}}\ldots,\mathbb{Z}_{k_{m_{3}}})$
=($={n_{1}-1\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}\choose m_{3}}\left(\frac{n_{1}U_{n}}{n_{1}-1}-\frac{V_{i,0,0}}{n_{1}-1}\right);$ (20)
(2) for $j=1,2,\dots,n_{2},$
$\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}\\ 1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{2}-1\\ j_{l}\neq j,\;l=1,2,\dots,m_{2}\\ 1\leq k_{1}<\ldots<k_{m_{3}}\leq n_{3}\end{subarray}}h(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},$ $\mathbb{Y}_{j_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}},\mathbb{Z}_{k_{1}}\ldots,\mathbb{Z}_{k_{m_{3}}})$
=($={n_{1}\choose m_{1}}{n_{2}-1\choose m_{2}}{n_{3}\choose m_{3}}\left(\frac{n_{2}U_{n}}{n_{2}-1}-\frac{V_{0,j,0}}{n_{2}-1}\right);$ (21)
(3) and for $k=1,2,\dots,n_{3},$
$\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}\\ 1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{2}\\ 1\leq k_{1}<\ldots<k_{m_{3}}\leq n_{3}-1\\ k_{l}\neq k,\;l=1,2,\dots,m_{3}\end{subarray}}h(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},$ $\mathbb{Y}_{j_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}},\mathbb{Z}_{k_{1}}\ldots,\mathbb{Z}_{k_{m_{3}}})$
=($={n_{1}\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}-1\choose m_{3}}\left(\frac{n_{3}U_{n}}{n_{3}-1}-\frac{V_{0,0,k}}{n_{3}-1}\right).$ (22)
Substituting (Appendix A: For Obtaining $EV_{l}$), (Appendix A: For Obtaining $EV_{l}$) and (Appendix A: For Obtaining $EV_{l}$) into (Appendix A: For Obtaining $EV_{l}$), and we obtain
$V_{l}=$ $nU_{n}-\frac{n(n-1)}{n-m}\frac{1}{{n_{1}\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}\choose m_{3}}}\Bigg{[}{n_{1}-1\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}\choose m_{3}}\left(\frac{n_{1}U_{n}}{n_{1}-1}-\frac{V_{l,0,0}}{n_{1}-1}\right)I_{[1,n_{1}]}(l)$
+($\qquad+{n_{1}\choose m_{1}}{n_{2}-1\choose m_{2}}{n_{3}\choose m_{3}}\left(\frac{n_{2}U_{n}}{n_{2}-1}-\frac{V_{0,l-n_{1},0}}{n_{2}-1}\right)I_{[n_{1}+1,n_{1}+n_{2}]}(l)$
+($\qquad+{n_{1}\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}-1\choose m_{3}}\left(\frac{n_{3}U_{n}}{n_{3}-1}-\frac{V_{0,0,l-n_{1}-n_{2}}}{n_{3}-1}\right)I_{[n_{1}+n_{2}+1,n]}(l)\Bigg{]}$
$=$ $nU_{n}-\frac{n(n-1)}{n-m}U_{n}\Bigg{[}\frac{n_{1}-m_{1}}{n_{1}(n_{1}-1)}I_{[1,n_{1}]}(l)+\frac{n_{2}-m_{2}}{n_{2}(n_{2}-1)}I_{[n_{1}+1,n_{1}+n_{2}]}(l)$
$\qquad+\frac{n_{3}-m_{3}}{n_{3}(n_{3}-1)}I_{[n_{1}+n_{2}+1,n]}(l)\Bigg{]}+\Bigg{[}\frac{n_{1}-m_{1}}{n_{1}(n_{1}-1)}V_{l,0,0}I_{[1,n_{1}]}(l)$
$\qquad+\frac{n_{2}-m_{2}}{n_{2}(n_{2}-1)}V_{0,l-n_{1},0}I_{[n_{1}+1,n_{1}+n_{2}]}(l)+\frac{n_{3}-m_{3}}{n_{3}(n_{3}-1)}V_{0,0,l-n_{1}-n_{2}}I_{[n_{1}+n_{2}+1,n]}(l)\Bigg{]}.$ (23)
Therefore the possible values of $EV_{l}$ are as follows
$EV_{l}=$ $n\theta-\frac{n(n-1)}{n-m}\theta\Bigg{[}\frac{n_{1}-m_{1}}{n_{1}(n_{1}-1)}I_{[1,n_{1}]}(l)+\frac{n_{2}-m_{2}}{n_{2}(n_{2}-1)}I_{[n_{1}+1,n_{1}+n_{2}]}(l)+\frac{n_{3}-m_{3}}{n_{3}(n_{3}-1)}I_{[n_{1}+n_{2}+1,n]}(l)\Bigg{]}$
$\qquad+\Bigg{[}\frac{n_{1}-m_{1}}{n_{1}(n_{1}-1)}\theta I_{[1,n_{1}]}(l)+\frac{n_{2}-m_{2}}{n_{2}(n_{2}-1)}\theta I_{[n_{1}+1,n_{1}+n_{2}]}(l)+\frac{n_{3}-m_{3}}{n_{3}(n_{3}-1)}\theta I_{[n_{1}+n_{2}+1,n]}(l)\Bigg{]}.$
Appendix B: Proof of Theorem 2.2
To prove the Theorem 2.2, we first provide some lemmas, which play a vital role in the proof.
The following lemma ensures both the existence and uniqueness of the solution to equation (16). Its proof is analogous to the one given by [7] for the two-sample case, hence omitted.
Lemma 6.1.
Under conditions (a) and (b) of the Theorem 2.2, we have
$P\left\{\min_{1\leq i\leq n}\left(V_{i}-EV_{i}\right)<0<\max_{1\leq i\leq n}\left(V_{i}-EV_{i}\right)\right\}\longrightarrow 1\;\;\;\text{as }\;\min\{n_{1},n_{2},n_{3}\}\rightarrow\infty.$
Lemma 6.2.
Suppose that conditions (a) and (b) of the Theorem 2.2 hold and let $S_{n}^{2}=\frac{1}{n}\sum_{l=1}^{n}\left(V_{l}-EV_{l}\right)^{2}$. Then,
$S_{n}^{2}=nS_{n_{1},n_{2},n_{3}}^{2}+o_{p}(1)\;\;\;\;\text{as }\;\min\{n_{1},n_{2},n_{3}\}\rightarrow\infty.$
Proof.
Using (23), for $1\leq l\leq n_{1}$, we have
$V_{l}=$ $\left(n-\frac{n(n-1)(n_{1}-m_{1})}{(n-m)(n_{1}-1)}\right)U_{n}+\frac{n(n-1)(n_{1}-m_{1})}{n_{1}(n-m)(n_{1}-1)}V_{l,0,0}$
Therefore,
$V_{l}-EV_{l}=$ $\frac{n(n-1)(n_{1}-m_{1})}{n_{1}(n-m)(n_{1}-1)}\left(V_{l,0,0}-U_{n}\right)+\frac{n}{n-m}\left(n-m-\frac{(n-1)(n_{1}-m_{1})}{n_{1}}\right)\left(U_{n}-\theta\right)$
and
$\frac{1}{n_{1}}\sum_{l=1}^{n_{1}}\left(V_{l,0,0}-U_{n}\right)\left(U_{n}-\theta\right)=\left(U_{n}-\theta\right)\left(\frac{1}{n_{1}}\sum_{l=1}^{n_{1}}V_{l,0,0}-U_{n}\right)=0.$
So, we have that
$\sum_{l=1}^{n_{1}}(V_{l}-EV_{l})^{2}=$ $\left[\frac{n(n-1)(n_{1}-m_{1})}{n_{1}(n-m)(n_{1}-1)}\right]^{2}\sum_{l=1}^{n_{1}}\left(V_{l,0,0}-U_{n}\right)^{2}$
$\qquad+\left[\frac{n}{n-m}\left(n-m-\frac{(n-1)(n_{1}-m_{1})}{n_{1}}\right)\right]^{2}n_{1}\left(U_{n}-\theta\right)^{2}.$
Similarly, for $n_{1}<l\leq n_{1}+n_{2}$, we have
$\sum_{l=n_{1}+1}^{n_{1}+n_{2}}(V_{l}-EV_{l})^{2}=$ $\left[\frac{n(n-1)(n_{2}-m_{2})}{n_{2}(n-m)(n_{2}-1)}\right]^{2}\sum_{l=n_{1}+1}^{n_{1}+n_{2}}\left(V_{0,l,0}-U_{n}\right)^{2}$
$\quad+\left[\frac{n}{n-m}\left(n-m-\frac{(n-1)(n_{2}-m_{2})}{n_{2}}\right)\right]^{2}n_{2}\left(U_{n}-\theta\right)^{2}.$
And similarly, for $n_{1}+n_{2}<l\leq n$, we have
$\sum_{l=n_{1}+n_{2}+1}^{n}(V_{l}-EV_{l})^{2}=$ $\left[\frac{n(n-1)(n_{3}-m_{3})}{n_{3}(n-m)(n_{3}-1)}\right]^{2}\sum_{l=n_{1}+n_{2}+1}^{n}\left(V_{0,0,l}-U_{n}\right)^{2}$
$\qquad+\left[\frac{n}{n-m}\left(n-m-\frac{(n-1)(n_{3}-m_{3})}{n_{3}}\right)\right]^{2}n_{3}\left(U_{n}-\theta\right)^{2}.$
Therefore,
$S_{n}^{2}=$ $\frac{1}{n}\sum_{l=1}^{n}\left(V_{l}-EV_{l}\right)^{2}$
$=$ $\frac{1}{n}\left[\frac{n(n-1)}{n-m}\right]^{2}\bigg{[}\frac{(n_{1}-m_{1})^{2}}{n_{1}(n_{1}-1)}\frac{1}{n_{1}(n_{1}-1)}\sum_{l=1}^{n_{1}}\left(V_{l,0,0}-\bar{V}_{\cdot,0,0}\right)^{2}$
$+\frac{(n_{2}-m_{2})^{2}}{n_{2}(n_{2}-1)}\frac{1}{n_{2}(n_{2}-1)}\sum_{l=n_{1}+1}^{n_{1}+n_{2}}\left(V_{0,l,0}-\bar{V}_{0,\cdot,0}\right)^{2}+\frac{(n_{3}-m_{3})^{2}}{n_{3}(n_{3}-1)}\frac{1}{n_{3}(n_{3}-1)}\left(V_{0,0,l}-\bar{V}_{0,0,\cdot}\right)^{2}\bigg{]}$
$+\frac{1}{n}\left[\frac{n}{n-m}\right]^{2}\bigg{[}n_{1}\left(n-m-\frac{(n-1)(n_{1}-m_{1})}{n_{1}}\right)^{2}+n_{2}\left(n-m-n_{2}\frac{(n-2)(n_{2}-m_{2})}{n_{2}}\right)^{2}$
$\qquad\qquad\qquad\qquad+n_{3}\left(n-m-\frac{(n-1)(n_{3}-m_{3})}{n_{3}}\right)^{2}\bigg{]}\left(U_{n}-\theta\right)^{2}.$ (*)
From part (a) of Theorem 2.1, we have $U_{n}-\theta=O_{p}\left(\min\{n_{1},n_{2},n_{3}\}^{-1/2}\right)$. Therefore, the second term in Equation (*) is
$\frac{1}{n}\left[\frac{n}{n-m}\right]^{2}\bigg{[}n_{1}\left(n-m-\frac{(n-1)(n_{1}-m_{1})}{n_{1}}\right)^{2}+n_{2}\left(n-m-\frac{(n-1)(n_{2}-m_{2})}{n_{2}}\right)^{2}$
$\qquad\qquad\qquad+n_{3}\left(n-m-\frac{(n-1)(n_{3}-m_{3})}{n_{3}}\right)^{2}\bigg{]}\left(U_{n}-\theta\right)^{2}=O_{p}(\min\{n_{1},n_{2},n_{3}\}^{-1}).$
Moreover, the first term of equation (*) is (see Chapter 2 of [25])
$\frac{1}{n}\left[\frac{n(n-1)}{n-m}\right]^{2}\bigg{[}\frac{(n_{1}-m_{1})^{2}}{n_{1}(n_{1}-1)}\frac{1}{n_{1}(n_{1}-1)}\sum_{l=1}^{n_{1}}\left(V_{l,0,0}-\bar{V}_{\cdot,0,0}\right)^{2}$
$\quad+\frac{(n_{2}-m_{2})^{2}}{n_{2}(n_{2}-1)}\frac{1}{n_{2}(n_{2}-1)}\sum_{l=n_{1}+1}^{n_{1}+n_{2}}\left(V_{0,l,0}-\bar{V}_{0,\cdot,0}\right)^{2}+\frac{(n_{3}-m_{3})^{2}}{n_{3}(n_{3}-1)}\frac{1}{n_{3}(n_{3}-1)}\left(V_{0,0,l}-\bar{V}_{0,0,\cdot}\right)^{2}\bigg{]}$
$=n\hat{\sigma}^{2}+o_{p}(1).$
By substituting the above into equation (*) and using Theorem 2.1, we obtain
$S_{n}^{2}=nS_{n_{1},n_{2},n_{3}}^{2}+o_{p}(1).$
This prove the assertion. ∎
The proof of the following Lemma is straight forward and on the same lines as of [1], and hence omitted.
Lemma 6.3.
Suppose that conditions (a) and (b) of the Theorem 2.2 hold. Then,
$\max_{1\leq l\leq n}\left|V_{l}-EV_{l}\right|=o_{p}\left(n^{1/2}\right)\;\;\text{ and }\;\;\frac{1}{n}\sum_{l=1}^{n}\left|V_{l}-EV_{l}\right|^{3}=o_{p}\left(n^{1/2}\right).$
Lemma 6.4.
Under the conditions (a) and (b) of the Theorem 2.2, we have
$|\lambda|=O_{p}\left(n^{-1/2}\right)\;\;\text{ and }\;\;\lambda=\frac{\left(U-\theta\right)}{S_{n}^{2}}+o_{p}\left(n^{-1/2}\right).$
Proof.
Noting that, (16) together with the fundamental inequality $\mid x\pm$ $y|\geq|x|-|y\mid$ leads to
00 0 $=\left|\frac{1}{n}\sum_{i=1}^{n}\frac{V_{i}-EV_{i}}{1+\lambda\left(V_{i}-EV_{i}\right)}\right|=\frac{1}{n}\left|\sum_{i=1}^{n}\left(V_{i}-EV_{i}\right)-\lambda\sum_{i=1}^{n}\frac{\left(V_{i}-EV_{i}\right)^{2}}{1+\lambda\left(V_{i}-EV_{i}\right)}\right|$
$\geq\frac{1}{n}\left|\lambda\sum_{i=1}^{n}\frac{\left(V_{i}-EV_{i}\right)^{2}}{1+\lambda\left(V_{i}-EV_{i}\right)}\right|-\frac{1}{n}\left|\sum_{i=1}^{n}\left(V_{i}-EV_{i}\right)\right|$
$\geq\frac{|\lambda|S_{n}^{2}}{1+|\lambda|\left(\max_{1\leq l\leq n}\left|V_{l}-EV_{l}\right|\right)}-\left|U_{n}-\theta\right|$
Using the Theorem 2.1, the second term is $O_{p}\left(n_{1}^{-1/2}\right)$ and, from Lemma 6.2, we obtain
$|\lambda|\left(1+|\lambda|\left(\max_{1\leq l\leq n}\left|V_{l}-EV_{l}\right|\right)\right)^{-1}=O_{p}\left(n^{-1/2}\right).$
Therefore, using $\max_{1\leq l\leq n}\left|V_{l}-EV_{l}\right|=o_{p}\left(n^{1/2}\right)$ from Lemma 6.3, we have
$|\lambda|=O_{p}\left(n^{-1/2}\right).$ (24)
To prove that second part of lemma, let $\gamma_{l}=\lambda\left(V_{l}-EV_{l}\right),\;l=1,2,\dots,n$. Then, using Lemma 6.3, we get
$\max_{1\leq l\leq n}\left|\gamma_{l}\right|$ $=|\lambda|\max_{1\leq l\leq n}\left|V_{l}-EV_{l}\right|=O_{p}\left(n^{-1/2}\right)o\left(n^{1/2}\right)=o_{p}(1).$ (25)
From (16), we have
00 0 $=\frac{1}{n}\sum_{i=1}^{n}\frac{V_{i}-EV_{i}}{1+\lambda\left(V_{i}-EV_{i}\right)}$
$=\frac{1}{n}\sum_{i=1}^{n}\left(V_{i}-EV_{i}\right)-\lambda\frac{1}{n}\sum_{i=1}^{n}\frac{\left(V_{i}-EV_{i}\right)^{2}}{1+\lambda\left(V_{i}-EV_{i}\right)}$
$=\frac{1}{n}\sum_{i=1}^{n}\left(V_{i}-EV_{i}\right)-\lambda\frac{1}{n}\sum_{i=1}^{n}\frac{\left(\left(V_{i}-EV_{i}\right)^{2}(1+\gamma_{i})-\lambda\left(V_{i}-EV_{i}\right)^{3}\right)}{1+\gamma_{i}}$
$=U_{n}-\theta-\lambda S_{n}^{2}+\lambda^{2}\frac{1}{n}\sum_{i=1}^{n}\frac{\left(V_{i}-EV_{i}\right)^{3}}{1+\gamma_{i}}.$
Also, using Lemma 6.3 and equations (24)-(25), last term in above expression is bounded by
$\frac{\frac{1}{n}\sum_{i=1}^{n}\left|V_{i}-EV_{i}\right|^{3}}{\left|1+\max_{1\leq l\leq n}\left|\gamma_{l}\right|\right|}\lambda^{2}=o_{p}\left(n^{1/2}\right)o_{p}(1)O_{p}\left(n^{-1}\right)=o_{p}\left(n^{-1/2}\right)$
Therefore, we get
$\lambda=\frac{\left(U-\theta\right)}{S_{n}^{2}}+o_{p}\left(n^{-1/2}\right).$
Hence the lemma is proved. ∎
Proof.
of Theorem 2.2: Now, we are ready to prove the theorem. Take $\gamma_{l}=\lambda\left(V_{l}-EV_{l}\right),\;l=1,2,\dots,n,$ and consider the $\log$ likelihood ratio given by
$-2\log R(\theta)$ $=2\sum_{1=1}^{n}\log\left(1+\lambda\left(V_{i}-EV_{i}\right)\right)$
$=2\sum_{1=1}^{n}\log\left(1+\gamma_{i}\right).$
From the Taylor series expansion for $\log\left(1+\gamma_{i}\right)$, we obtain
$-2\log R(\theta)=2\sum_{1=1}^{n}\gamma_{i}-\sum_{1=1}^{n}{\gamma_{i}}^{2}+\frac{2}{3}\sum_{1=1}^{n}\gamma_{i}{}^{3}-\frac{1}{2}\sum_{1=1}^{n}{\gamma_{i}}^{4}+\ldots$ (26)
Let us first consider the third term of (26)
$\frac{2}{3}\sum_{1=1}^{n}\gamma_{i}{}^{3}=\frac{2}{3}\sum_{1=1}^{n}\lambda^{3}\left(V_{i}-EV_{i}\right)^{3}$ $\leq\frac{2}{3}\,n^{\frac{3}{2}}|\lambda|^{3}\frac{1}{n}\sum_{1=1}^{n}\left|V_{i}-EV_{i}\right|^{3}\frac{1}{\sqrt{n}}$
$=n^{\frac{3}{2}}O_{p}(n^{-3/2})o_{p}(n^{\frac{1}{2}})o_{p}(n^{-\frac{1}{2}})$ (By using Lemmas 6.3 and 6.4)
$=o_{p}(1).$
Hence, it is straightforward to verify that all terms with powers of $\gamma_{i}$ equal to or greater than 3 will be $o_{p}(1)$.
Next, we consider the second term of (26) as
$\sum_{1=1}^{n}\gamma_{i}^{2}=\sum_{1=1}^{n}\lambda^{2}\left(V_{i}-\theta\right)^{2}=n\lambda^{2}S_{n}^{2}$ $\leq n|\lambda|^{2}(nS_{n_{1},n_{2},n_{3}}^{2}+o_{p}(1))$ (By using Lemma 6.2)
$=n\;o_{p}(n^{-1})(nS_{n_{1},n_{2},n_{3}}^{2}+o_{p}(1))$ (By using Lemma 6.4)
$=o_{p}(1).$
Now, consider the first term of (26)
$2\sum_{1=1}^{n}\gamma_{i}=2\sum_{1=1}^{n}\lambda\left(V_{i}-EV_{i}\right)$ $=2\lambda n\frac{1}{n}\sum_{1=1}^{n}\left(V_{i}-EV_{i}\right)$
$=2\lambda n(U_{n}-\theta)$
$=\frac{n(U_{n}-\theta)^{2}}{S_{n}^{2}}+no_{p}(n^{-1/2})(U_{n}-\theta)$ (By using Lemma 6.4)
$=\frac{n(U_{n}-\theta)^{2}}{S_{n}^{2}}+n\,o_{p}(n^{-1/2})O_{p}(n^{-1/2})$ (By using Theorem 2.1)
$=\frac{n(U_{n}-\theta)^{2}}{S_{n}^{2}}+o_{p}(1).$
Therefore,
$-2\log R(\theta)=\frac{n(U_{n}-\theta)^{2}}{S_{n}^{2}}+o_{p}(1)=\frac{(U_{n}-\theta)^{2}}{S_{n_{1},n_{2},n_{3}}^{2}}+o_{p}(1)\;\;\;\;\text{as }\;n_{1}\rightarrow\infty.$ (By using Lemma 6.2)
By using part (b) of Theorem 2.1 and the Slutsky’s theorem, $-2\log R(\theta)$ converges in distribution to a $\chi^{2}$ random variable with one degree of freedom. Hence the theorem proved.