### v2

# Jackknife Empirical Likelihood for Multivariate three-Sample U-Statistics and its Applications

Abstract.
We develop a jackknife empirical likelihood (JEL) framework for inference on parameters defined through multivariate three-sample U-statistic. From three independent multivariate samples, we construct JEL ratio statistic based on suitable jackknife pseudo-values and, under mild regularity conditions, establish a Wilks-type result showing that the log JEL ratio converges in distribution to a chi-square limit. This provides asymptotically valid confidence intervals for the parameter of interest without explicit variance estimation or heavy resampling. To illustrate the usefulness of the proposed method, we construct confidence intervals for differences in volume under the surface (VUS) measures, which are widely used in classification problems. Through Monte Carlo simulations, we compare the performance of JEL-based confidence intervals with those obtained from normal approximation of U-statistic and kernel-based methods. The findings indicate that the proposed JEL approach outperforms existing methods in terms of coverage probability and computational efficiency. Finally, we apply our methods to a recent real dataset.
Keyword: Jackknife empirical likelihood (JEL); Volume under surface (VUS); U-statistics.
$\dagger$}Corresponding Author mail: skkattu@isichennai.res.in
1. Introduction
Empirical likelihood (EL) has emerged as a useful tool in statistical inference when traditional parametric assumptions about the distribution of the data are either unknown or do not hold. EL operates on the principle of maximizing a nonparametric likelihood function with constraints on observed data with applications to hypotheses testing and confidence interval estimation. Early work on EL is due to Owen (owen1988empirical, owen1990empirical). owen2001empirical highlighted that empirical likelihood (EL) inherently constructs confidence regions, accommodates supplementary constraints, and preserves the natural scale and transformations of the data. This adaptability has established EL as a useful methodology in in handling incomplete data and addressing complex multivariate problems (jian2008).
For nonlinear statistics such as U-statistics, however, the original EL formulation can be difficult to implement directly, because the constraints involve nonlinear functionals of the empirical distribution. To address this difficulty, jing2009jackknife proposed the jackknife empirical likelihood (JEL) method, which constructs empirical likelihood ratios based on jackknife pseudo-values. Their work established Wilks-type results for one- and two-sample univariate U-statistics and demonstrated that JEL often improves coverage accuracy over normal approximation methods while avoiding explicit variance estimation. Since then, JEL has been extended to various settings, including multivariate mean comparisons Wang2013, non-smooth estimating equations Li2016, and tests based on energy statistics SANG2020, underscoring its flexibility for complex nonparametric functionals. sang2021jackknife developed a jackknife empirical likelihood procedure for nonparametric $k$-sample tests based on categorical Gini correlations, and established a chi-square limiting distribution for the resulting test statistic.
Despite this progress, existing JEL developments for U-statistics are predominantly limited to one- or two-sample problems, or to vector parameters in a single-sample setting. In many applications, however, the parameter of interest is naturally expressed as a functional of multivariate three-sample U-statistics. Examples include ordering probabilities in three-arm clinical trials, multivariate stress–strength reliability problems, and measures of three-class separation in diagnostic and classification settings. In such problems, the underlying U-statistic is built from three independent multivariate samples, and the usual one- or two-sample JEL theory is not directly applicable. A general JEL framework for multivariate $3$-sample U-statistics is therefore needed.
The goal of this paper is to develop such a framework. We consider U-statistics built from three independent $q$-dimensional samples with degrees $(m_{1},m_{2},m_{3})$ and a symmetric kernel, and define a combined-sample construction that allows us to work with jackknife pseudo-values that are asymptotically independent but not identically distributed. Within this setting, we obtain a jackknife empirical likelihood ratio for a general scalar parameter $\theta$ defined through a three-sample U-statistic and establish a Wilks-type theorem: under mild moment and sample-size conditions, the JEL log-likelihood ratio converges in distribution to a chi-square random variable with one degree of freedom. This yields asymptotically valid confidence intervals for $\theta$ without explicit variance estimation or resampling.
To illustrate the practical relevance of the proposed methodology, we focus on problems arising in diagnostic accuracy assessment. For three-class classification, we construct confidence intervals for the difference between two correlated volumes under ROC surfaces (VUSs). In this case, the parameters of interest can be written as functionals of multivariate three-sample U-statistics, and our general JEL theory provides a coherent nonparametric inference procedure. Extensive Monte Carlo simulations under several multivariate models, including the Marshall–Olkin bivariate exponential and FGM-based bivariate Pareto distributions, are used to study the finite-sample performance of the proposed JEL intervals in comparison with intervals based on asymptotic normality with jackknife variance estimation and kernel-based bootstrap methods. Finally, we apply the methods to biomarker data from the Alzheimer’s Disease Neuroimaging Initiative, where we construct confidence intervals for differences in VUSs when comparing diagnostic markers across cognitive groups.
The paper is structured as follows. Section 2 we study $q$-dimensional U-statistics based on three samples. We study its asymptotic properties and find a consistent estimator of the unknown variance of the U-statistic based on Jackknife pseudo values. Jackknife empirical likelihood approach for these U-statistics is also discussed. As an illustration,in Section 3, we find confidence intervals for the difference between two dependent volumes under surfaces (VUS) using asymptotic normality of U-statistics, JEL method and bootstrapping of kernel type estimators . Section 4 presents extensive simulation studies assessing the performance of these confidence intervals under the Marshall–Olkin bivariate exponential and Bivariate Pareto models. Section 4 illustrates the methodology through a real-data analysis. Section 5 highlights the findings of the study along with some open problems.
2. Three sample multivariate U-Statistics
Suppose $\left(\mathbb{X}_{1},\ldots,\mathbb{X}_{n_{1}}\right),\left(\mathbb{Y}_{1},\ldots,\mathbb{Y}_{n_{2}}\right)$ and $\left(\mathbb{Z}_{1},\ldots,\mathbb{Z}_{n_{3}}\right)$ represents three independent random samples from $q$-dimensional distribution functions $F$, $G$ and $H$, respectively, where $\mathbb{X}_{i}=(X_{i1},X_{i2},\dots,X_{iq})^{T},\;i=1,2,...,n_{1}$, $\mathbb{Y}_{j}=(Y_{j1},Y_{j2},\dots,Y_{jq})^{T},\;$ $j=1,2,...,n_{2},$ and $\mathbb{Z}_{k}=(Z_{k1},Z_{k2},\dots,Z_{kq})^{T},\;k=1,2,...,n_{3}$, and $q\geq 1$.
Let
$\theta=E[h\left(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},\mathbb{Y}_{j_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}},\mathbb{Z}_{k_{1}}\ldots,\mathbb{Z}_{k_{m_{3}}}\right)]\in\Re$
be the parameteric function of interest, where $h:\mathcal{X}^{m_{1}}\times\mathcal{Y}^{m_{2}}\times\mathcal{Z}^{m_{3}}\;\rightarrow\;\Re$ is a measurable function symmetric in its arguments, i.e., invariant under any permutation of indices within each sample, and $\mathcal{X},\mathcal{Y},$ and $\mathcal{Z}$ be the sample spaces associated with the distribution functions $F$, $G$, and $H$, respectively.
An unbiased estimator of $\theta$ based on three-sample U-statistic with a symmetric kernel $h$ of degree $\left(m_{1},m_{2},m_{3}\right)$ is given by
$U_{n_{1},n_{2},n_{3}}=$ $\frac{1}{{n_{1}\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}\choose m_{3}}}\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}\\ 1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{2}\\ 1\leq k_{1}<\ldots<k_{m_{3}}\leq n_{3}\end{subarray}}h\left(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},\mathbb{Y}_{j_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}},\mathbb{Z}_{k_{1}}\ldots,\mathbb{Z}_{k_{m_{3}}}\right),$ (2.1)
where $\{i_{1},\ldots,i_{m_{1}}\}$, $\{j_{1},\ldots,j_{m_{2}}\}$ and $\{k_{1},\ldots,k_{m_{3}}\}$ indicates $m_{1}$, $m_{2}$ and $m_{3}$ integers chosen without replacement from sets $\{1,\ldots,n_{1}\}$, $\{1,\ldots,n_{2}\}$ and $\{1,\ldots,n_{3}\}$, respectively.
2.1. Three sample multivariate U-Statistic -Asymptotic results
In this section we study the asymptotic properties of the U-statistic $U_{n_{1},n_{2},n_{3}}$. Define
$\sigma_{100}^{2}$ $=Var\left[E\left(h\left(\mathbb{X}_{1},\dots,\mathbb{X}_{m_{1}},\mathbb{Y}_{1},\dots,\mathbb{Y}_{m_{2}},\mathbb{Z}_{1},\dots,\mathbb{Z}_{m_{3}}\right)\mid\mathbb{X}_{1}\right)\right],$
$\sigma_{010}^{2}$ $=Var\left[E\left(h\left(\mathbb{X}_{1},\dots,\mathbb{X}_{m_{1}},\mathbb{Y}_{1},\dots,\mathbb{Y}_{m_{2}},\mathbb{Z}_{1},\dots,\mathbb{Z}_{m_{3}}\right)\mid\mathbb{Y}_{1}\right)\right]$
and
$\sigma_{001}^{2}$ $=Var\left[E\left(h\left(\mathbb{X}_{1},\dots,\mathbb{X}_{m_{1}},\mathbb{Y}_{1},\dots,\mathbb{Y}_{m_{2}},\mathbb{Z}_{1},\dots,\mathbb{Z}_{m_{3}}\right)\mid\mathbb{Z}_{1}\right)\right].$
Theorem 2.1.
(a) $U_{n_{1},n_{2},n_{3}}\xrightarrow{\text{ a.s }}\theta$ as $\min\{n_{1},n_{2},n_{3}\}\rightarrow\infty$;
(b) Suppose that $\sigma_{100}^{2},\sigma_{010}^{2},\sigma_{001}^{2}>0$ holds. Let $\sigma^{2}=\frac{m_{1}^{2}\sigma_{100}^{2}}{n_{1}}+\frac{m_{2}^{2}\sigma_{010}^{2}}{n_{2}}+\frac{m_{3}^{2}\sigma_{001}^{2}}{n_{3}},$ Then,
$\frac{U_{n_{1},n_{2},n_{3}}-\theta}{\sigma}\xrightarrow{d}N(0,1),\text{ as }\min\{n_{1},n_{2},n_{3}\}\rightarrow\infty.$
For the proof of part (a) of the above theorem, see pp. 151–153 of korolyuk. The proof of part (b), i.e., the asymptotic normality of $U_{n_{1},n_{2},n_{3}}$, follows from puri1971nonparametric.
Since $\sigma^{2}$ is unknown, we find its consistent estimator. Following steps outline how to define a jackknife pseudo-value-based estimator. Consider
(1) the U statistics based on all observations $U_{n_{1},n_{2},n_{3}}$, as defined by (2.1);
(2) for $i=1,2,\dots,n_{1},$ the U statistics after deleting observation $X_{i}$, is given by
$U_{n_{1}-1,n_{2},n_{3}}^{(-i)}=\frac{1}{{n_{1}-1\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}\choose m_{3}}}\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}-1\\ i_{l}\neq i,\;l=1,2,\dots,m_{1},\\ 1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{2}\\ 1\leq k_{1}<\ldots<k_{m_{3}}\leq n_{3}\end{subarray}}\thinspace h\left(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},\mathbb{Y}_{j_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}},\mathbb{Z}_{k_{1}}\ldots,\mathbb{Z}_{k_{m_{3}}}\right);$ (2.2)
(3) for $j=1,2,\dots,n_{2},$ the U statistics after deleting observation $Y_{j}$, is given by
$U_{n_{1},n_{2}-1,n_{3}}^{(-j)}=\frac{1}{{n_{1}\choose m_{1}}{n_{2}-1\choose m_{2}}{n_{3}\choose m_{3}}}\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}\\ 1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{2}-1\\ j_{l}\neq j,\;l=1,2,\dots,m_{2},\\ 1\leq k_{1}<\ldots<k_{m_{3}}\leq n_{3}\end{subarray}}\thinspace h\left(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},\mathbb{Y}_{j_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}},\mathbb{Z}_{k_{1}}\ldots,\mathbb{Z}_{k_{m_{3}}}\right);$ (2.3)
(4) for $k=1,2,\dots,n_{3},$ the U statistics after deleting observation $Z_{k}$, is given by
$U_{n_{1},n_{2},n_{3}-1}^{(-k)}=\frac{1}{{n_{1}\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}-1\choose m_{3}}}\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}\\ 1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{2}\\ 1\leq k_{1}<\ldots<k_{m_{3}}\leq n_{3}-1\\ k_{l}\neq k,\;l=1,2,\dots,m_{3}\end{subarray}}\thinspace h\left(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},\mathbb{Y}_{j_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}},\mathbb{Z}_{k_{1}}\ldots,\mathbb{Z}_{k_{m_{3}}}\right).$ (2.4)
The jackknife pseudo-values for individual samples are defined as,
$V_{i,0,0}=n_{1}U_{n_{1},n_{2},n_{3}}-\left(n_{1}-1\right)U_{n_{1}-1,n_{2},n_{3}}^{(-i)},\;\;\forall\;\;i=1,2,\dots,n_{1},$ (2.5)
$V_{0,j,0}=n_{2}U_{n_{1},n_{2},n_{3}}-\left(n_{2}-1\right)U_{n_{1},n_{2}-1,n_{3}}^{(-j)},\;\;\forall\;\;j=1,2,\dots,n_{2}$ (2.6)
and
$V_{0,0,k}=n_{3}U_{n_{1},n_{2},n_{3}}-\left(n_{3}-1\right)U_{n_{1},n_{2},n_{3}-1}^{(-k)},\;\;\forall\;\;k=1,2,\dots,n_{3}.$ (2.7)
It is easy to verify that $\bar{V}_{\cdot,0,0}=\bar{V}_{0,\cdot,0}=\bar{V}_{0,0,\cdot}=U_{n_{1},n_{2},n_{3}}$, where
$\bar{V}_{\cdot,0,0}=\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}V_{i,0,0},\;\;\bar{V}_{0,\cdot,0}=\frac{1}{n_{2}}\sum_{j=1}^{n_{2}}V_{0,j,0}\;\text{ and }\;\bar{V}_{0,0,\cdot}=\frac{1}{n_{3}}\sum_{k=1}^{n_{3}}V_{0,0,k}.$
The jackknife pseudo-values based estimator of $\sigma^{2}$, proposed by arvesen1969jackknifing (also see wang2010empirical, guangming2013nonparametric and an2018jackknife),is given as
$\hat{\sigma}^{2}=:\frac{1}{n_{1}\left(n_{1}-1\right)}\sum_{i=1}^{n_{1}}\left(V_{i,0,0}-\bar{V}_{\cdot,0,0}\right)^{2}$ $+\frac{1}{n_{2}\left(n_{2}-1\right)}\sum_{j=1}^{n_{2}}\left(V_{0,j,0}-\bar{V}_{0,\cdot,0}\right)^{2}$
$+\frac{1}{n_{3}\left(n_{3}-1\right)}\sum_{k=1}^{n_{3}}\left(V_{0,0,k}-\bar{V}_{0,0,\cdot}\right)^{2},$ (2.8)
where $V_{i,0,0}$, $V_{0,j,0}$ and $V_{0,0,k}$ are defined in equations (2.5), (2.6) and (2.7), respectively.
Lemma 2.1.
Suppose that $\sigma_{100}^{2},\sigma_{010}^{2},\sigma_{001}^{2}>0$. Then
$\hat{\sigma}^{2}-\sigma^{2}=o_{p}\!\left(\min\{n_{1},n_{2},n_{3}\}^{-1}\right),$
where $\sigma^{2}=\frac{m_{1}^{2}\sigma_{100}^{2}}{n_{1}}+\frac{m_{2}^{2}\sigma_{010}^{2}}{n_{2}}+\frac{m_{3}^{2}\sigma_{001}^{2}}{n_{3}},$ and $\hat{\sigma}^{2}$ is defined in (2.1).
Proof: The proof can be easily derived using Chebyshev’s inequality and is therefore omitted (see guangming2013nonparametric).
Remark 2.2.
(i) From Theorem 2.1 we get that $U_{n_{1},n_{2},n_{3}}$ is a consistent estimator of $\theta$.
(ii) From Lemma 2.1 we get that $\hat{\sigma}^{2}$ is a consistent estimator of $\sigma^{2}$.
(iii) From Theorem 2.1 and Lemma 2.1,it follows that a two-sided $(1-\alpha)$ level asymptotic CI for $\theta$ is given by
[…]
where $z_{\alpha}$ is the upper $\alpha$-th percentile point of a standard normal distribution.
2.2. Three sample multivariate U-Statistics - JEL
jing2009jackknife showed that the JEL method, based on one and two-sample U-statistics, outperforms the normal approximation method. We study generalization of the JEL method to accommodate q-dimensional U-statistics arising from three samples. By broadening the applicability of JEL beyond conventional one and two-sample univariate settings, our approach enhances its relevance to vector valued three-sample statistical data analysis.
We combine the three samples and treat them as a single sample. For this purpose, let $n=n_{1}+n_{2}+n_{3}$ and define $\mathbb{W}_{l}$ as
$\mathbb{W}_{l}=\begin{cases}\mathbb{X}_{l},&\text{ if }l=1,\dots,n_{1}\\ \mathbb{Y}_{l-n_{1}},&\text{ if }l=n_{1}+1,\dots,n_{1}+n_{2}\\ \mathbb{Z}_{l-n_{1}-n_{2}},&\text{ if }l=n_{1}+n_{2}+1,\dots,n\end{cases}.$
Therefore, $(\mathbb{W}_{1},\mathbb{W}_{2},\dots,\mathbb{W}_{n})=(\mathbb{X}_{1},\dots,\mathbb{X}_{n_{1}},\mathbb{Y}_{1},\dots,\mathbb{Y}_{n_{2}},\mathbb{Z}_{1},\dots,\mathbb{Z}_{n_{3}})$, where $\mathbb{W}_{i}^{\prime}s$ are independent but need not be identically distributed. Then, the U-statistic based on single sample $(\mathbb{W}_{1},\mathbb{W}_{2},\dots,\mathbb{W}_{n})$ with degree $m=m_{1}+m_{2}+m_{3}$ is given as follows:
$U_{n}=\frac{1}{{n\choose m}}\mathop{\sum}_{\begin{subarray}{c}1\leq l_{1}<\ldots<l_{m}\leq n\end{subarray}}\tilde{h}\left(\mathbb{W}_{l_{1}},\mathbb{W}_{l_{2}},\ldots,\mathbb{W}_{l_{m}}\right),$ (2.9)
where
$\tilde{h}\left(\mathbb{W}_{l_{1}},\mathbb{W}_{l_{2}},\ldots,\mathbb{W}_{l_{m}}\right)$
$\begin{cases}&\text{if}\;\;1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}\\ c^{*}\;h\left(\mathbb{W}_{i_{1}},\dots,\mathbb{W}_{i_{m_{1}}},\mathbb{W}_{j_{1}},\ldots,\mathbb{W}_{j_{m_{2}}},\mathbb{W}_{k_{1}},\dots,\mathbb{W}_{k_{m_{3}}}\right),&n_{1}+1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{1}+n_{2}\\ &n_{1}+n_{2}+1\leq k_{1}<\ldots<k_{m_{3}}\leq n\\ 0,&\text{otherwise}\end{cases}$
with $c^{*}=\frac{{n\choose m}}{{n_{1}\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}\choose m_{3}}}$. Note that $U_{n_{1},n_{2},n_{3}}=U_{n}$.
Consider the U-statistic after deleting observation $\mathbb{W}_{l}$, from sample $(\mathbb{W}_{1},\mathbb{W}_{2},\dots,\mathbb{W}_{n})$, given by
$U_{n-1}^{(-l)}=\frac{1}{{n-1\choose m}}\mathop{\sum}_{\begin{subarray}{c}1\leq l_{1}<\ldots<l_{m}\leq n\\ l_{i}\neq l,\;i=1,2,\dots,m\end{subarray}}\tilde{h}\left(\mathbb{W}_{l_{1}},\mathbb{W}_{l_{2}},\ldots,\mathbb{W}_{l_{m}}\right),\;\;\forall\;l=1,2,\dots,n.$ (2.10)
Define the jackknife pseudo-values $V_{l}$, for combined sample $(\mathbb{W}_{1},\mathbb{W}_{2},\dots,\mathbb{W}_{n})$ as
$V_{l}=nU_{n}-(n-1)U_{n-1}^{(-l)},\;\;\forall\;l=1,2,\dots,n.$ (2.11)
It is easy to verify that $U_{n}=\frac{1}{n}\sum_{i=1}^{n}V_{i},$ where $V_{i}$ ’s are asymptotically independent random variables (shi1984approximate).
The expected values of $V_{l}^{\prime}s$ are given below (details are provided in Appendix A).
$EV_{l}=\begin{cases}\theta\left(\frac{n}{n-m}\right)\left[n-m-\frac{(n-1)(n_{1}-m_{1})}{n_{1}}\right],&\text{ if }\;l=1,2,\ldots,n_{1}\\ \penalty 10000\ \\ \theta\left(\frac{n}{n-m}\right)\left[n-m-\frac{(n-1)(n_{2}-m_{2})}{n_{2}}\right],&\text{ if }\;l=n_{1}+1,n_{1}+2,\ldots,n_{1}+n_{2}\\ \penalty 10000\ \\ \theta\left(\frac{n}{n-m}\right)\left[n-m-\frac{(n-1)(n_{3}-m_{3})}{n_{3}}\right],&\text{ if }\;l=n_{1}+n_{2}+1,\ldots,n\end{cases}.$ (2.12)
It is easy to verify that $\frac{1}{n}\sum_{l=1}^{n}EV_{l}=\theta$ and, in particular, if $m_{1}=m_{2}=m_{3}$ and $n_{1}=n_{2}=n_{3}$, we get $EV_{l}=\theta$, $\forall\;l=1,2,\ldots,n$.
Finally, we employ the EL method to construct the confidence intervals for $\theta$ by maximizing the empirical likelihood function subject to the constraints based on jackknife pseudo values $V_{l}^{\prime}$s. Let $\mathbf{p}=\left(p_{1},\ldots,p_{n}\right)$ be a probability vector assigning probability $p_{i}$ to each $V_{i}$. The JEL evaluated at $\theta$ becomes
$L(\theta)=\max\left\{\prod_{i=1}^{n}p_{i};\sum_{i=1}^{n}p_{i}=1,\sum_{i=1}^{n}p_{i}\left(V_{i}-EV_{i}\right)=0\right\}.$ (2.13)
We know that $\prod_{i=1}^{n}p_{i}$ subject to $\sum_{i=1}^{n}p_{i}=1$ attain its maximum at $\frac{1}{n^{n}}$ at $p_{i}=\frac{1}{n}$. Hence the jackknife empirical likelihood ratio for $\theta$ is given by
$R(\theta)=\frac{L(\theta)}{n^{-n}}=\max\left\{\prod_{i=1}^{n}\left(np_{i}\right);\sum_{i=1}^{n}p_{i}=1,\sum_{i=1}^{n}p_{i}\left(V_{i}-EV_{i}\right)=0\right\}.$ (2.14)
Using Lagrange multipliers method, when
$\min_{1\leq i\leq n}\left(V_{i}-EV_{i}\right)<0<\max_{1\leq i\leq n}\left(V_{i}-EV_{i}\right),$
we obtain
$p_{i}=\frac{1}{n}\frac{1}{1+\lambda\left(V_{i}-EV_{i}\right)},\;\;\forall\;\;i=1,2,\dots,n,$ (2.15)
where $\lambda$ satisfies
$\frac{1}{n}\sum_{i=1}^{n}\frac{V_{i}-EV_{i}}{1+\lambda\left(V_{i}-EV_{i}\right)}=0.$ (2.16)
Hence we obtain the jackknife empirical log-likelihood ratio as
$\log R(\theta)=-\sum_{i=1}^{n}\log\left(1+\lambda\left(V_{i}-EV_{i}\right)\right).$
We can use $\log R(\theta)$ for constructing confidence intervals or developing test on $\theta$. For this, we need to obtain the asymptotic distribution of $\log R(\theta)$. We next derive the limiting distribution of the jackknife empirical log-likelihood ratio as an analog to Wilk’s theorem. The proof of the following theorem is given in Appendix B.
Theorem 2.3.
Suppose that following conditions holds:
(a) $E\left(h^{2}\left(\mathbb{X}_{1},\ldots,\mathbb{X}_{m_{1}},\mathbb{Y}_{1},\ldots,\mathbb{Y}_{m_{2}},\mathbb{Z}_{1}\ldots,\mathbb{Z}_{m_{3}}\right)\right)<\infty$ and $\sigma_{100}^{2},\sigma_{010}^{2},\sigma_{001}^{2}>0$,
(b) $0<\lim\inf n_{1}/n_{2}\leq\lim\sup n_{1}/n_{2}<\infty$ and $0<\lim\inf n_{2}/n_{3}\leq\lim\sup n_{2}/n_{3}<\infty$.
Then, as $n\rightarrow\infty,-2\log R(\theta)$ converges in distribution to a $\chi^{2}$ random variable with one degree of freedom.
3. Application: Volumes under ROC surfaces
The receiver operating characteristic (ROC) curve and its associated area under the curve (AUC) are fundamental tools for assessing the diagnostic accuracy of binary classifiers, summarizing their ability to discriminate between two outcome states across all possible threshold values. In medical science, ROC analysis is routinely used to evaluate diagnostic tests, compare biomarkers, and determine optimal thresholds for disease detection (yang2024transformed).
However, in medical diagnostics, disease severity may exist on a continuum—such as healthy, intermediate, and diseased — rather than fitting neatly into a binary framework. For example, consider levels of three biomarkers used to assess the risk of developing cardiovascular disease: let $X$ be the level of Biomarker 1 (e.g., LDL cholesterol), $Y$ be the level of Biomarker 2 (e.g., HDL cholesterol), and $Z$ be the level of Biomarker 3 (e.g., triglycerides). We are interested in finding the probability that a patient has an LDL cholesterol level $X$ that is lower than their HDL cholesterol level $Y$, and their HDL cholesterol level $Y$ that is lower than their triglyceride level $Z$, i.e., $P(X<Y<Z)$. This probability, is called Volume under ROC surface and denoted by VUS. It can be interpreted as a measure of how often the biomarker levels are ordered in a specific way that might be associated with different risk levels or health conditions (he2008).
This extension takes care of the inherent limitations of binary classifications in medical diagnostics, where conditions often manifest along a spectrum rather than as discrete categories. tian2011exact. wang2010empirical, guangming2013nonparametric, and an2018jackknife considered non-parametric inference for VUS using the JEL method.
Suppose $X_{i}=\left(X_{i1},X_{i2}\right)^{T},i=1,2,\ldots,n_{1},Y_{j}=\left(Y_{j1},Y_{j2}\right)^{T},j=1,2,\ldots,n_{2}$, and $Z_{k}=\left(Z_{k1},Z_{k2}\right)^{T},k=1,2,\ldots,n_{3},$ represent random samples of three populations having independent joint distribution functions $F,G$ and $H$, respectively.
The difference of two correlated VUS’s can be defined as
$\theta_{V}=P\left(X_{11}<Y_{11}<Z_{11}\right)-P\left(X_{12}<Y_{12}<Z_{12}\right).$
In the following part of this subsection, we use three different non-parametric methods for finding confidence intervals for $\theta_{V}$. We adopt the notations introduced in Sections 2.1 and 2.2, with $q=2$ and $m_{1}=m_{2}=m_{3}=1$.
U-statistic based confidence intervals for $\theta_{V}$:
We can rewrite $\theta_{V}$ as
$\theta_{V}$ $=E\left[I\left(X_{11}<Y_{11}<Z_{11}\right)-I\left(X_{12}<Y_{12}<Z_{12}\right)\right],$
where $I\left(x<y<z\right)=\begin{cases}1,&\text{if }\;x<y<z\\ 0,&\text{ otherwise}\end{cases}$.
Consider the symmetric kernel $h^{VUS}$ given by
$h^{VUS}\left(\left(X_{i1},X_{i2}\right),\left(Y_{j1},Y_{j2}\right),\left(Z_{k1},Z_{k2}\right)\right)=I\left(X_{i1}<Y_{j1}<Z_{k1}\right)-I\left(X_{i2}<Y_{j2}<Z_{k2}\right).$
An estimator of $\theta_{V}$ can be found as U-statistic with kernel $h$ as
$U_{n_{1},n_{2},n_{3}}^{VUS}=\frac{1}{n_{1}n_{2}n_{3}}\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}\sum_{k=1}^{n_{3}}h^{VUS}\left(\left(X_{i1},X_{i2}\right),\left(Y_{j1},Y_{j2}\right),\left(Z_{k1},Z_{k2}\right)\right).$
The U-statistic $U_{n_{1},n_{2},n_{3}}^{VUS}$ is an unbiased and consistent estimator of $\theta$ (Lehmann, (1951)). Also, using Theorem 2.1, a two sided asymptotic $(1-\alpha)$ level CI for $\theta_{V}$ based on the asymptotic distribution of $U_{n_{1},n_{2},n_{3}}^{VUS}$ is
[…]
where $\hat{\sigma}$ is the jackknife pseudo-values based consistent estimator of $Var(U_{n_{1},n_{2},n_{3}}^{VUS})$ (see arvesen1969jackknifing), as defined by (2.1). Note that this confidence interval was also considered in an2018jackknife.
JEL based confidence interval for $\theta_{V}$:
We use Theorem 2.3 to find JEL based confidence intervals for $\theta_{V}$. We can define $\log R(\theta_{V})$ on the same lines as of Section 2.2. Hence, the JEL based confidence interval for $\theta_{V}$ at level $(1-\alpha)$ is given by
$CI=\{\theta_{V}\;|\;-2\log R(\theta_{V})\leq\chi^{2}_{1,1-\alpha}\}.$ (3.2)
Note that JEL based confidence interval was also considered in an2018jackknife.
Kernel based confidence interval for $\theta_{V}$: To estimate the difference of two VUS’s kang2013 proposed the kernel-smoothed version of the difference of two VUS’s as follows;
${\hat{\theta}_{V}}=\frac{1}{n_{1}n_{2}n_{3}}\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}\sum_{k=1}^{n_{3}}\left(\Phi\left(\frac{Y_{1j}-X_{1i}}{\sqrt{b_{1}^{2}+b_{2}^{2}}}\right)\Phi\left(\frac{Z_{1k}-Y_{1j}}{\sqrt{b_{2}^{2}+b_{3}^{2}}}\right)-\Phi\left(\frac{Y_{2j}-X_{2i}}{\sqrt{b_{4}^{2}+b_{5}^{2}}}\right)\Phi\left(\frac{Z_{2k}-Y_{2j}}{\sqrt{b_{5}^{2}+b_{6}^{2}}}\right)\right),$ (3.3)
where $\Phi(\cdot)$ represents the standard normal distribution function and the bandwidths are calculated by
$b_{i}=0.9\min\left\{sd\left(x_{i}\right)\right.,\left.\operatorname{iqr}\left(x_{i}\right)/1.34\right\}n_{i}^{-0.2},\;\;\;\;i=1,2,\dots,6,$
where $sd(\cdot)$ and $\operatorname{iqr}(\cdot)$ represent the standard deviation and the interquartile range. silverman2018 introduced the bandwidth $b_{i}$ which controls the level of smoothing in kernel estimation. This asymptotic bandwidth selection method has proven effective across a wide range of density functions and has been widely adopted by researchers in diagnostic studies.
We use the estimator (3.3) and bootstrap method to obtain a kernel-based confidence interval for $\theta_{V}$. In each bootstrap replication, we compute a point estimate $\hat{\theta}_{V}$ of $\theta_{V}$, which allows us to construct the confidence interval using the percentile method. The kernel-based bootstrap confidence interval for $\theta_{V}$ is obtained in Step 5 of Algorithm 1. This approach is commonly used for constructing confidence intervals; see polansky2001bandwidth, feinberg2024kernel, and lin2025bootstrap.
Algorithm 1 Bootstrap Percentile CI for a Kernel-Based Estimator of $\theta_{V}$
1:Compute the original estimate $\hat{\theta}_{V}$ (as defined by (3.3)).
2:for $b=1,\dots,B$ do
3:a: Draw bootstrap samples $(X_{1}^{*(b)},\dots,X_{n_{1}}^{*(b)})$, $(Y_{1}^{*(b)},\dots,Y_{n_{2}}^{*(b)})$ and $(Z_{1}^{*(b)},\dots,Z_{n_{3}}^{*(b)})$ by sampling with replacement from $\{X_{1},\dots,X_{n_{1}}\}$, $\{Y_{1},\dots,Y_{n_{2}}\}$ and $\{Z_{1},\dots,Z_{n_{3}}\}$, respectively.
4:b: Based on these bootstrap samples, compute the bootstrap estimate $\hat{\theta}_{V}^{*(b)}$ using (3.3).
5:Sort $\hat{\theta}_{V}^{*(1)},\dots,\hat{\theta}_{V}^{*(B)}$ in ascending order to obtain
$\hat{\theta}^{*(1)}_{V}\leq\cdots\leq\hat{\theta}^{*(B)}_{V}$.
6:Let
$\hat{\theta}^{*}_{\alpha/2}=\hat{\theta}^{*(\lceil B(\alpha/2)\rceil)}_{V},\quad\hat{\theta}^{*}_{1-\alpha/2}=\hat{\theta}^{*(\lceil B(1-\alpha/2)\rceil)}_{V}$,
where $\lceil x\rceil$ is the smallest integer greater than or equal to $x$.
7:The bootstrap percentile $(1-\alpha)$-confidence interval for $\theta_{V}$ is
$\bigl[\hat{\theta}^{*}_{\alpha/2},\,\hat{\theta}^{*}_{1-\alpha/2}\bigr].$ (3.4)
Another possible application of the methodology discussed in Section 2 is the assessment of a diagnostic test based on imaging features. Each patient is represented by a feature vector $\mathbb{x}\in\Re^{q}$, and the diagnostic procedure assigns a real-valued score via a function $s:\Re^{q}\to\Re$. Patients are to be classified into three ordered categories (e.g., healthy, mildly diseased, severely diseased) on the basis of this score.
Let $\mathbb{X}$, $\mathbb{Y}$, and $\mathbb{Z}$ denote independent observations from the healthy, mildly diseased, and severely diseased populations, respectively. The parameter of interest is
$\theta_{s}=P\big(s(\mathbb{X})<s(\mathbb{Y})<s(\mathbb{Z})\big),$
which is the probability that a randomly selected triplet, one patient from each group, is correctly ordered by the scoring function $s(\cdot)$. A natural point estimator of $\theta_{s}$ provides a summary measure of the accuracy of the diagnostic test, and one may also be interested in constructing a confidence interval for $\theta_{s}$.
4. Simulation
In this section, we employ Monte Carlo simulation studies to evaluate the finite sample performance of CIs for difference of two VUS’s. The three methods used are the JEL, normal approximation with jackknife variance estimator and kernel based bootstrapping. We assess the performance of each method using two distinct criteria:
(a) Coverage Probability (CP): This represents the probability that the true parameter value falls within the constructed confidence interval (CI). Methods with lower coverage error—defined as the discrepancy between the actual coverage probability and the nominal value are preferred.
(b) Average Length (AL) of CIs: Methods that produce CIs with shorter average lengths are desirable, as shorter intervals provide more precise information about the location of the unknown parameter.
We conduct simulation studies with different sample sizes and simulation is repeated 1,000 time. We use R package “empilik” for finding JEL based confidence intervals. For constructing confidence intervals based on kernel based method we use 1000 bootstrap replications. Two bivariate probability models are considered for the simulation studies.
Marshall-Olkin bivariate exponential model (MOBVE($\lambda_{1},\lambda_{2},\lambda_{3}$)) : We consider the Marshall-Olkin bivariate exponential model MOBVE($\lambda_{1},\lambda_{2},\lambda_{3}$), introduced by marshall1967multivariate. The joint cumulative distribution function of this bivariate model is given by
$F_{X_{1},X_{2}}(x_{1},x_{2})=1-\exp[-\lambda_{1}x_{1}-\lambda_{2}x_{2}-\lambda_{3}\max\{x_{1},x_{2}\}],\;\;x_{1}>0,\;x_{2}>0,$
where $\lambda_{i}>0,\;i=1,2,3.$ We generated sample sizes $n_{1}$, $n_{2}$, and $n_{3}$ from three independent models MOBVE($\lambda_{x_{1}},\lambda_{x_{2}},\lambda_{x_{3}}$), MOBVE($\lambda_{y_{1}},\lambda_{y_{2}},\lambda_{y_{3}}$), and MOBVE($\lambda_{z_{1}},\lambda_{z_{2}},\lambda_{z_{3}}$), respectively.
In Table 2, we explored various configurations for sample sizes $(n_{1},n_{2},n_{3})$ and the parameters $(\lambda_{x_{1}},\lambda_{x_{2}},\lambda_{x_{3}};\lambda_{y_{1}},\lambda_{y_{2}},\lambda_{y_{3}};\lambda_{z_{1}},\lambda_{z_{2}},\lambda_{z_{3}})$. This comprehensive investigation aims to evaluate the performance and efficacy of these three methods under various parameter settings within the MOBVE framework.
Bivariate Pareto model : In this part, we explore the Farlie-Gumbel-Morgenstern (FGM) copula based on two independent Pareto distributions with C.D.F.s defined as follows:
$F_{X_{i}}(x_{i})=1-\left(\frac{\lambda_{i}}{x_{i}}\right)^{\alpha_{i}},\;x_{i}\geq\lambda_{i}>0,\;\alpha_{i}>0,\;i=1,2.$
For $1\leq\theta\leq 1$, the FGM copula, or joint C.D.F., is given by
$C_{\theta}(F_{X_{1}}(x_{1}),F_{X_{2}}(x_{2}))$ $=$ $F_{X_{1}}(x_{1})F_{X_{2}}(x_{2})[1+\theta(1-F_{X_{1}}(x_{1}))(1-F_{X_{2}}(x_{2}))],$
$\qquad\qquad\qquad 0\leq F_{X_{1}}(x_{1}),\,F_{X_{2}}(x_{2})\leq 1.$
We refer to this bivariate model as the FGM-based Bivariate Pareto model (FGMPM), denoted as FGMPM($\lambda_{1},\lambda_{2},\alpha_{1},\alpha_{2},\theta$). For simulations, we generated sample sizes $n_{1}$, $n_{2}$, and $n_{3}$ from three FGMPM models: $FGMPM(\lambda_{x_{1}},\lambda_{x_{2}},\alpha_{x_{1}},\alpha_{x_{2}},\theta_{x}),\;FGMPM(\lambda_{y_{1}},\lambda_{y_{2}},\alpha_{y_{1}},\alpha_{y_{2}},\theta_{y}),$ and $FGMPM(\lambda_{z_{1}},\lambda_{z_{2}},\alpha_{z_{1}},\alpha_{z_{2}},\theta_{z}),$ respectively. The samples were generated using the “rCopula” function in the “copula” package in R. Results of the simulation studies for this model are presented in Table 3 for various configurations of sample sizes $(n_{1},n_{2},n_{3})$ and parameters $(\lambda_{x_{1}},\lambda_{x_{2}},\alpha_{x_{1}},\alpha_{x_{2}},\theta_{x};\lambda_{y_{1}},\lambda_{y_{2}},\alpha_{y_{1}},\alpha_{y_{2}},\theta_{y};\lambda_{z_{1}},\lambda_{z_{2}},\alpha_{z_{1}},\alpha_{z_{2}},\theta_{z})$.
The simulations were conducted on a laptop with the following hardware configuration: Processor: Intel(R) Core(TM) i7-7500U CPU @ 2.70GHz, 4 cores; Memory: 8 GB DDR4 RAM; Storage: 240 GB SSD WD green; Operating System: Windows 10, 64-bit.
The approximate time taken by the three methods for obtaining results during simulations is presented in Table 1.
The following observations can be made from Tables 1- 3:
1. We observe that for various configurations of parameters of the Marshall-Olkin bivariate exponential and bivariate Pareto models, the average lengths (ALs) of the confidence intervals based on JEL, normal approximation, and kernel-based methods decrease as the sample sizes increase.
2. In most cases, as the sample size increases, the coverage probabilities of the confidence intervals generated by the three methods tend to align with the nominal level (95%). This is anticipated because larger sample sizes yield more detailed data insights.
3. From Tables 2-3, we can see that CI based on JEL method achieves closer to 95% coverage across various choices of parameters compared to the other two methods. The average interval lengths are generally better for CIs based on the normal approximation and kernel-based methods for small sample sizes. However, the average lengths are almost same for moderate sample sizes.
4. From Table 1, we can observe that the simulation time taken using the normal approximation with the jackknife variance estimator and kernel-based bootstrapping methods is significantly higher than the time taken by the JEL method.
Therefore, based on the simulation studies, we recommend the JEL method as it requires significantly less time to construct confidence intervals compared to the other two methods and also provides better coverage probability.
Note that an2018jackknife also conducted simulation studies for the confidence intervals defined in (3.1) and (3.2). However, they did not consider the bivariate Pareto model and did not compare these methods with kernel-based confidence intervals.
Table 1. Simulation time taken by various methods under two models
Model
Samples Size
$(n_{1},n_{2},n_{3})$
JEL Normal Approx Kernel
MOBVE $(10,10,10)$ 1.6 minutes 1.95 minutes 24 minutes
$(100,100,100)$ 40 hours 69 hours 240 hours
bivariate Pareto $(10,10,10)$ 2.63 minutes 4.6 minutes 50 minutes
$(100,100,100)$ 135 hours 246 hours 656 hours
Table 2. Coverage probability (CP) and average length (AL) of confidence intervals for the difference two VUS’s at 95% confidence level: Under MOBVE models
$(\lambda_{x_{1}},\lambda_{x_{2}},\lambda_{x_{3}};\lambda_{y_{1}},\lambda_{y_{2}},\lambda_{y_{3}};\lambda_{z_{1}},\lambda_{z_{2}},\lambda_{z_{3}})$ $(n_{1},n_{2},n_{3})\;\;\;$ JEL Normal Approx Kernel CP (%) AL CP (%) AL CP (%) AL (1,1,1;1,1,1;1,1,1) (10,10,10) 96.8 .501 97.8 0.384 97 0.392 (20,20,20) 96.5 .332 96.8 0.325 95.8 0.341 (30,30,30) 95.7 .229 97 0.257 96.6 0.264 (50,50,50) 95 .18 96.6 0.183 96 0.195 (70,70,70) 94.4 .126 94 0.124 93.5 0.129 (100,100,100) 95.8 .102 94.2 0.103 94 0.122 (1,2,0;1,1,0;2,1,0) (10,10,10) 97 .564 92.8 0.516 91.8 0.504 (20,20,20) 96.4 .35 93 0.361 92 0.341 (30,30,30) 95.8 .248 92.5 0.239 93 0.25 (50,50,50) 95.4 .187 94 0.187 92.5 0.169 (70,70,70) 95.2 .157 94 0.187 92.5 0.158 (100,100,100) 95.2 .131 94.5 0.130 94 0.135 (1/3,1/3,2/3;2/3,2/3,4/3;1,1,2) (10,10,10) 99.4 .263 99.6 0.234 97.5 0.21 (20,20,20) 97 .169 97.5 0.151 98.5 0.161 (30,30,30) 96 .112 97.5 0.121 98.25 0.125 (50,50,50) 95.6 .085 95.75 0.096 97.5 0.089 (70,70,70) 95.6 .067 94.5 0.069 96.5 0.071 (100,100,100) 95.2 .053 94.8 0.058 96 0.062 (3/5,3/5,2/5;6/5,6/5,4/5;9/5,9/5,6/5) (10,10,10) 99.2 .319 98.75 0.269 96.5 0.235 (20,20,20) 95.2 .2 97 0.197 97.25 0.211 (30,30,30) 97 .146 96.5 0.151 97 0.142 (50,50,50) 96.4 .103 96.5 0.098 97 0.106 (70,70,70) 96 .087 96 0.082 96.5 0.09 (100,100,100) 95.6 .066 96 0.069 96 0.065 (1/19,1/19,18/19;2/19,2/19,36/19;3/19,3/19,54/19) (10,10,10) 98 0.09 86 0.07 88 0.079 (20,20,20) 95.2 0.071 98 0.06 98 0.055 (30,30,30) 97 0.048 99 0.052 94 0.049 (50,50,50) 91 0.034 89 0.034 88.5 0.036 (70,70,70) 90.4 0.027 90 0.028 88 0.029 (100,100,100) 93.6 0.023 91 0.022 90 0.025
Table 3. Coverage probability (CP) and average length (AL) of confidence intervals for the difference two VUS’s at 95% confidence level: Under Bivariate Pareto models
$(\lambda_{x_{1}},\lambda_{x_{2}},\alpha_{x_{1}},\alpha_{x_{2}},\theta_{x};\lambda_{y_{1}},\lambda_{y_{2}},\alpha_{y_{1}},\alpha_{y_{2}},\theta_{y};\lambda_{z_{1}},\lambda_{z_{2}},\alpha_{z_{1}},\alpha_{z_{2}},\theta_{z})$ $(n_{1},n_{2},n_{3})\;\;\;$ JEL Normal Approx Kernel CP (%) AL CP (%) AL CP (%) AL (1,1,1,1,0.5;1,1,1,1,0.5;1,1,1,1,-0.5) (10,10,10) 96 .493 96.3 0.451 98 0.485 (20,20,20) 97 .311 97 0.298 96.6 0.322 (30,30,30) 95 .274 94.3 0.262 94.1 0.281 (50,50,50) 96.8 .192 93.8 0.201 93.7 0.182 (70,70,70) 95.2 .156 94.8 0.153 94 0.159 (100,100,100) 95.4 .109 94.8 0.110 94.2 0.101 (1,2,1,2,0.5;2,1,2,1,0.5;2,2,1,1,-0.5) (10,10,10) 96 .557 91 0.549 89 0.514 (20,20,20) 98 .362 99 0.38 91 0.391 (30,30,30) 92 .268 92 0.306 91 0.321 (50,50,50) 98 .193 99 0.235 93 0.242 (70,70,70) 97 .198 94 0.189 92.5 0.191 (100,100,100) 96 .183 94 0.185 94 0.187 (5,1,1,5,0.2;5,1,2,2,0.5;1,5,5,1,0.9) (10,10,10) 97 .460 90 0.465 88 0.45 (20,20,20) 93 .345 92 0.325 91 0.34 (30,30,30) 94 .275 92 0.295 92 0.29 (50,50,50) 96 .174 97 0.17 98 0.183 (70,70,70) 96 .165 93.5 0.166 94 0.169 (100,100,100) 95.7 .151 94 0.150 94 0.153 (0.5,1,0.5,1,-0.5;1,1,0.5,0.5,-0.5;1,1,5,1,-0.2) (10,10,10) 94 .359 92 0.356 91 0.341 (20,20,20) 98 .212 93 0.201 92 0.204 (30,30,30) 97 .169 93 0.157 92 0.17 (50,50,50) 96 .127 93 0.13 92.5 0.129 (70,70,70) 95.8 .098 93.5 0.095 93 0.097 (100,100,100) 95.5 .076 94 0.075 93.5 0.076 (10,1,0.5,10,-0.9;1,1,15,0.2,-0.1;0.2,1,5,0.1,0.9) (10,10,10) 92 .530 90 0.521 89 0.532 (20,20,20) 98 .317 91 0.29 89 0.299 (30,30,30) 97 .262 91 0.267 91 0.276 (50,50,50) 96 .196 92 0.196 93 0.189 (70,70,70) 96 .163 98 0.165 99 0.17 (100,100,100) 95.8 .123 94 0.12 93.5 0.125
5. Data analysis
[…]
$\theta_{V}=P(X_{1}<Y_{1}<Z_{1})-P(X_{2}<Y_{2}<Z_{2}),$
[…]
In total, 883 results were included in this analysis, with sample sizes for the three individual classes being 222, 122, and 539, respectively. We used the JEL, Normal approximation, and Kernel-based methods, as described in Section 3, to construct confidence intervals for the difference between two VUSs. The JEL method took 3.4 hours, the Normal approximation method took 4.7 hours, and the Kernel-based method (with 1000 bootstraps) took 52 hours to obtain results for this real data analysis using R software on the previously specified laptop.
[…]
Note that, for the application, we consider a recent dataset where the biomarkers and disease stages are entirely distinct from those used in the real-data analysis in an2018jackknife.
Table 4. Confidence Intervals for $\theta$ at different confidence levels using different methods
Method JEL Normal Approx Kernel
Confidence Interval 90% (0.3124,0.3753) (0.3161,0.3723) (0.3083,0.3743)
Confidence Interval 95% (0.3062,0.3813) (0.3107,0.3776) (0.3006,0.3803)
Confidence Interval 99% (0.3007,0.3866) (0.3002,0.3882) (0.2909,0.3904)
6. Concluding remarks
In this paper, we studied U-statistics based on three multivariate samples and developed a jackknife empirical likelihood (JEL) methodology for such statistics. This would extend the applicability of JEL methods to three sample multivariate problems.
We constructed confidence intervals for differences in two dependent VUSs using three nonparametric procedures. We conducted an extensive simulation study under the Marshall-Olkin bivariate exponential and Bivariate Pareto distributions. Our results showed the superior performance of the JEL method in terms of coverage probability and computational complexity when compared to the normal approximation using jackknife pseudo-values based variance estimator and kernel-based methods.
Additionally, we conducted a real data analysis using a recent Alzheimer’s disease dataset to illustrate the practical benefits of the JEL method in real-world scenarios. This study highlights the superior performance, versatility, lower computational cost and practical benefits of the JEL method over its competitors recommending its use in statistical practices that require reliable nonparametric inference for complex data structures.
Beyond diagnostic testing, the proposed framework is also relevant in other scientific domains. For example, an ecologist could be interested in comparison of multiple species in three habitats: protected forest, logged forest and agricultural land. He would be interested in testing the hypothesis that the distribution of multiple species in these three habitats are identical. rizzo2016energy have discussed energy based statistics for this testing problem. Many of them can be expressed as U-statistics. Asymptotic distributions of the test statistics (either U-statistics or JEL ratio) will follow from the results in Section 2.
We are working on $q$-dimensional $k$-sample U-statistics and the asymptotic distribution of the associated JEL ratio. One possible application of the theory developed would be construction of confidence intervals for hypervolume under the ROC manifold which is a generalization of ROC curves to multi-category classification. The results will be presented elsewhere.
Appendix A: For obtaining $EV_{l}$
We simplify the expression for $V_{l}^{\prime}s$ and then find $EV_{l}$.
For simplicity, denote
$\bar{h}_{i_{1},\dots,i_{m_{1}},j_{1},\dots,j_{m_{2}},k_{1},\dots,k_{m_{3}}}\thinspace=h(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},\mathbb{Y}_{j_{1}-n_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}-n_{1}},\mathbb{Z}_{k_{1}-n_{1}-n_{2}}\ldots,\mathbb{Z}_{k_{m_{3}}-n_{1}-n_{2}}).$
For $l=1,2,\dots,n$, we have
$V_{l}=$ $nU_{n}-(n-1)U_{n-1}^{(-l)}$
$=$ $nU_{n}-\frac{(n-1)}{{n-1\choose m}}\mathop{\sum}_{\begin{subarray}{c}1\leq l_{1}<\ldots<l_{m}\leq n\\ l_{i}\neq l,\;i=1,2,\dots,m\end{subarray}}\tilde{h}\left(\mathbb{W}_{l_{1}},\mathbb{W}_{l_{2}},\ldots,\mathbb{W}_{l_{m}}\right)$
$=$ $nU_{n}-\frac{(n-1)}{{n-1\choose m}}\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\\ <j_{1}<\ldots<j_{m_{2}}\\ <k_{1}<\ldots<k_{m_{3}}\leq n-1\\ i_{i}\leq n_{1}<j_{j}\leq n_{1}+n_{2}<k_{k},\\ \forall\;i=1,\dots,m_{1},\;j=1,\dots,m_{2},\;k=1,\dots,m_{3}\\ i_{i}=j_{i}=k_{i}\neq l,\;i=1,2,\dots,m\end{subarray}}\frac{{n\choose m}}{{n_{1}\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}\choose m_{3}}}\bar{h}_{i_{1},\dots,i_{m_{1}},j_{1},\dots,j_{m_{2}},k_{1},\dots,k_{m_{3}}}$
$=$ $nU_{n}-\frac{n(n-1)}{n-m}\frac{1}{{n_{1}\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}\choose m_{3}}}\Bigg[\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}-1\\ i_{i}\neq l,\;i=1,2,\dots,m_{1}\\ n_{1}+1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{1}+n_{2}\\ n_{1}+n_{2}+1\leq k_{1}<\ldots<k_{m_{3}}\leq n\end{subarray}}\thinspace I_{[1,n_{1}]}(l)\,\bar{h}_{i_{1},\dots,i_{m_{1}},j_{1},\dots,j_{m_{2}},k_{1},\dots,k_{m_{3}}}$
$\qquad\qquad\qquad+\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}\\ n_{1}+1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{1}+n_{2}-1\\ j_{j}\neq l,\;j=1,2,\dots,m_{2}\\ n_{1}+n_{2}+1\leq k_{1}<\ldots<k_{m_{3}}\leq n\end{subarray}}I_{[n_{1}+1,n_{1}+n_{2}]}(l)\,\bar{h}_{i_{1},\dots,i_{m_{1}},j_{1},\dots,j_{m_{2}},k_{1},\dots,k_{m_{3}}}$
$\qquad\qquad\qquad+\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}\\ n_{1}+1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{1}+n_{2}\\ n_{1}+n_{2}+1\leq k_{1}<\ldots<k_{m_{3}}\leq n-1\\ k_{k}\neq l,\;k=1,2,\dots,m_{3}\end{subarray}}I_{[n_{1}+n_{2}+1,n]}(l)\,\bar{h}_{i_{1},\dots,i_{m_{1}},j_{1},\dots,j_{m_{2}},k_{1},\dots,k_{m_{3}}}\Bigg],$ (6.1)
where $I_{[a,b]}(x)$ is a indicator function, defined as $I_{[a,b]}(x)=\begin{cases}1,&\text{if }a\leq x\leq b\\ 0,&\text{otherwise}\end{cases}.$
Since $U_{n_{1},n_{2},n_{3}}=U_{n}$ and combining (2.2) and (2.5), (2.3) and (2.6), and (2.4) and (2.7), we have the following expressions,
(1) for $i=1,2,\dots,n_{1},$
$\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}-1\\ i_{l}\neq i,\;l=1,2,\dots,m_{1}\\ 1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{2}\\ 1\leq k_{1}<\ldots<k_{m_{3}}\leq n_{3}\end{subarray}}h(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},$ $\mathbb{Y}_{j_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}},\mathbb{Z}_{k_{1}}\ldots,\mathbb{Z}_{k_{m_{3}}})$
$={n_{1}-1\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}\choose m_{3}}\left(\frac{n_{1}U_{n}}{n_{1}-1}-\frac{V_{i,0,0}}{n_{1}-1}\right);$ (6.2)
(2) for $j=1,2,\dots,n_{2},$
$\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}\\ 1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{2}-1\\ j_{l}\neq j,\;l=1,2,\dots,m_{2}\\ 1\leq k_{1}<\ldots<k_{m_{3}}\leq n_{3}\end{subarray}}h(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},$ $\mathbb{Y}_{j_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}},\mathbb{Z}_{k_{1}}\ldots,\mathbb{Z}_{k_{m_{3}}})$
$={n_{1}\choose m_{1}}{n_{2}-1\choose m_{2}}{n_{3}\choose m_{3}}\left(\frac{n_{2}U_{n}}{n_{2}-1}-\frac{V_{0,j,0}}{n_{2}-1}\right);$ (6.3)
(3) and for $k=1,2,\dots,n_{3},$
$\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}\\ 1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{2}\\ 1\leq k_{1}<\ldots<k_{m_{3}}\leq n_{3}-1\\ k_{l}\neq k,\;l=1,2,\dots,m_{3}\end{subarray}}h(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},$ $\mathbb{Y}_{j_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}},\mathbb{Z}_{k_{1}}\ldots,\mathbb{Z}_{k_{m_{3}}})$
$={n_{1}\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}-1\choose m_{3}}\left(\frac{n_{3}U_{n}}{n_{3}-1}-\frac{V_{0,0,k}}{n_{3}-1}\right).$ (6.4)
Substituting (Appendix A: For obtaining $EV_{l}$), (Appendix A: For obtaining $EV_{l}$) and (Appendix A: For obtaining $EV_{l}$) into (Appendix A: For obtaining $EV_{l}$), and we obtain
$V_{l}=$ $nU_{n}-\frac{n(n-1)}{n-m}\frac{1}{{n_{1}\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}\choose m_{3}}}\Bigg[{n_{1}-1\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}\choose m_{3}}\left(\frac{n_{1}U_{n}}{n_{1}-1}-\frac{V_{l,0,0}}{n_{1}-1}\right)I_{[1,n_{1}]}(l)$
$\hskip 18.49988pt+{n_{1}\choose m_{1}}{n_{2}-1\choose m_{2}}{n_{3}\choose m_{3}}\left(\frac{n_{2}U_{n}}{n_{2}-1}-\frac{V_{0,l-n_{1},0}}{n_{2}-1}\right)I_{[n_{1}+1,n_{1}+n_{2}]}(l)$
$\hskip 18.49988pt+{n_{1}\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}-1\choose m_{3}}\left(\frac{n_{3}U_{n}}{n_{3}-1}-\frac{V_{0,0,l-n_{1}-n_{2}}}{n_{3}-1}\right)I_{[n_{1}+n_{2}+1,n]}(l)\Bigg]$
$=$ $nU_{n}-\frac{n(n-1)}{n-m}U_{n}\Bigg[\frac{n_{1}-m_{1}}{n_{1}(n_{1}-1)}I_{[1,n_{1}]}(l)+\frac{n_{2}-m_{2}}{n_{2}(n_{2}-1)}I_{[n_{1}+1,n_{1}+n_{2}]}(l)$
$\hskip 18.49988pt+\frac{n_{3}-m_{3}}{n_{3}(n_{3}-1)}I_{[n_{1}+n_{2}+1,n]}(l)\Bigg]+\Bigg[\frac{n_{1}-m_{1}}{n_{1}(n_{1}-1)}V_{l,0,0}I_{[1,n_{1}]}(l)$
$\hskip 18.49988pt+\frac{n_{2}-m_{2}}{n_{2}(n_{2}-1)}V_{0,l-n_{1},0}I_{[n_{1}+1,n_{1}+n_{2}]}(l)+\frac{n_{3}-m_{3}}{n_{3}(n_{3}-1)}V_{0,0,l-n_{1}-n_{2}}I_{[n_{1}+n_{2}+1,n]}(l)\Bigg].$ (6.5)
Therefore the values of $EV_{l}$ are as follows
$EV_{l}=$ $n\theta-\frac{n(n-1)}{n-m}\theta\Bigg[\frac{n_{1}-m_{1}}{n_{1}(n_{1}-1)}I_{[1,n_{1}]}(l)+\frac{n_{2}-m_{2}}{n_{2}(n_{2}-1)}I_{[n_{1}+1,n_{1}+n_{2}]}(l)+\frac{n_{3}-m_{3}}{n_{3}(n_{3}-1)}I_{[n_{1}+n_{2}+1,n]}(l)\Bigg]$
$\hskip 17.00024pt+\Bigg[\frac{n_{1}-m_{1}}{n_{1}(n_{1}-1)}\theta I_{[1,n_{1}]}(l)+\frac{n_{2}-m_{2}}{n_{2}(n_{2}-1)}\theta I_{[n_{1}+1,n_{1}+n_{2}]}(l)+\frac{n_{3}-m_{3}}{n_{3}(n_{3}-1)}\theta I_{[n_{1}+n_{2}+1,n]}(l)\Bigg].$
Appendix B: Proof of Theorem 2.3
We first provide some lemmas which play a vital role in the proof of Theorem 2.3.
The following lemma ensures both the existence and uniqueness of the solution to equation (2.16). Its proof is analogous to the one given by jing2009jackknife for the two-sample case, hence omitted.
Lemma 6.1.
Under conditions (a) and (b) of the Theorem 2.3, we have
$P\left\{\min_{1\leq i\leq n}\left(V_{i}-EV_{i}\right)<0<\max_{1\leq i\leq n}\left(V_{i}-EV_{i}\right)\right\}\longrightarrow 1\;\;\;\text{as }\;\min\{n_{1},n_{2},n_{3}\}\rightarrow\infty.$
Lemma 6.2.
Suppose that conditions (a) and (b) of the Theorem 2.3 hold and let $S_{n}^{2}=\frac{1}{n}\sum_{l=1}^{n}\left(V_{l}-EV_{l}\right)^{2}$. Then,
$S_{n}^{2}=n\sigma^{2}+o_{p}(1)\;\;\;\;\text{as }\;\min\{n_{1},n_{2},n_{3}\}\rightarrow\infty.$
Proof.
Using (6.5), for $1\leq l\leq n_{1}$, we have
$V_{l}=$ $\left(n-\frac{n(n-1)(n_{1}-m_{1})}{(n-m)(n_{1}-1)}\right)U_{n}+\frac{n(n-1)(n_{1}-m_{1})}{n_{1}(n-m)(n_{1}-1)}V_{l,0,0}$
Therefore,
$V_{l}-EV_{l}=$ $\frac{n(n-1)(n_{1}-m_{1})}{n_{1}(n-m)(n_{1}-1)}\left(V_{l,0,0}-U_{n}\right)+\frac{n}{n-m}\left(n-m-\frac{(n-1)(n_{1}-m_{1})}{n_{1}}\right)\left(U_{n}-\theta\right)$
and
$\frac{1}{n_{1}}\sum_{l=1}^{n_{1}}\left(V_{l,0,0}-U_{n}\right)\left(U_{n}-\theta\right)=\left(U_{n}-\theta\right)\left(\frac{1}{n_{1}}\sum_{l=1}^{n_{1}}V_{l,0,0}-U_{n}\right)=0.$
Hence, we have that
$\sum_{l=1}^{n_{1}}(V_{l}-EV_{l})^{2}=$ $\left[\frac{n(n-1)(n_{1}-m_{1})}{n_{1}(n-m)(n_{1}-1)}\right]^{2}\sum_{l=1}^{n_{1}}\left(V_{l,0,0}-U_{n}\right)^{2}$
$\qquad+\left[\frac{n}{n-m}\left(n-m-\frac{(n-1)(n_{1}-m_{1})}{n_{1}}\right)\right]^{2}n_{1}\left(U_{n}-\theta\right)^{2}.$
Similarly, for $n_{1}<l\leq n_{1}+n_{2}$, we have
$\sum_{l=n_{1}+1}^{n_{1}+n_{2}}(V_{l}-EV_{l})^{2}=$ $\left[\frac{n(n-1)(n_{2}-m_{2})}{n_{2}(n-m)(n_{2}-1)}\right]^{2}\sum_{l=n_{1}+1}^{n_{1}+n_{2}}\left(V_{0,l,0}-U_{n}\right)^{2}$
$\quad+\left[\frac{n}{n-m}\left(n-m-\frac{(n-1)(n_{2}-m_{2})}{n_{2}}\right)\right]^{2}n_{2}\left(U_{n}-\theta\right)^{2}.$
And for $n_{1}+n_{2}<l\leq n$, we have
$\sum_{l=n_{1}+n_{2}+1}^{n}(V_{l}-EV_{l})^{2}=$ $\left[\frac{n(n-1)(n_{3}-m_{3})}{n_{3}(n-m)(n_{3}-1)}\right]^{2}\sum_{l=n_{1}+n_{2}+1}^{n}\left(V_{0,0,l}-U_{n}\right)^{2}$
$\qquad+\left[\frac{n}{n-m}\left(n-m-\frac{(n-1)(n_{3}-m_{3})}{n_{3}}\right)\right]^{2}n_{3}\left(U_{n}-\theta\right)^{2}.$
Therefore,
$S_{n}^{2}=$ $\frac{1}{n}\sum_{l=1}^{n}\left(V_{l}-EV_{l}\right)^{2}$
$=$ $\frac{1}{n}\left[\frac{n(n-1)}{n-m}\right]^{2}\bigg[\frac{(n_{1}-m_{1})^{2}}{n_{1}(n_{1}-1)}\frac{1}{n_{1}(n_{1}-1)}\sum_{l=1}^{n_{1}}\left(V_{l,0,0}-\bar{V}_{\cdot,0,0}\right)^{2}$
$+\frac{(n_{2}-m_{2})^{2}}{n_{2}(n_{2}-1)}\frac{1}{n_{2}(n_{2}-1)}\sum_{l=n_{1}+1}^{n_{1}+n_{2}}\left(V_{0,l,0}-\bar{V}_{0,\cdot,0}\right)^{2}+\frac{(n_{3}-m_{3})^{2}}{n_{3}(n_{3}-1)}\frac{1}{n_{3}(n_{3}-1)}\left(V_{0,0,l}-\bar{V}_{0,0,\cdot}\right)^{2}\bigg]$
$+\frac{1}{n}\left[\frac{n}{n-m}\right]^{2}\bigg[n_{1}\left(n-m-\frac{(n-1)(n_{1}-m_{1})}{n_{1}}\right)^{2}+n_{2}\left(n-m-n_{2}\frac{(n-2)(n_{2}-m_{2})}{n_{2}}\right)^{2}$
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt+n_{3}\left(n-m-\frac{(n-1)(n_{3}-m_{3})}{n_{3}}\right)^{2}\bigg]\left(U_{n}-\theta\right)^{2}.$ (*)
From part (a) of Theorem 2.1, we have $U_{n}-\theta=O_{p}\left(\min\{n_{1},n_{2},n_{3}\}^{-1/2}\right)$. Therefore, the second term in Equation (*) is
$\frac{1}{n}\left[\frac{n}{n-m}\right]^{2}\bigg[n_{1}\left(n-m-\frac{(n-1)(n_{1}-m_{1})}{n_{1}}\right)^{2}+n_{2}\left(n-m-\frac{(n-1)(n_{2}-m_{2})}{n_{2}}\right)^{2}$
$\qquad\qquad\qquad+n_{3}\left(n-m-\frac{(n-1)(n_{3}-m_{3})}{n_{3}}\right)^{2}\bigg]\left(U_{n}-\theta\right)^{2}=O_{p}(\min\{n_{1},n_{2},n_{3}\}^{-1}).$
Moreover, the first term of equation (*) is (see Chapter 2 of wang2010empirical)
$\frac{1}{n}\left[\frac{n(n-1)}{n-m}\right]^{2}\bigg[\frac{(n_{1}-m_{1})^{2}}{n_{1}(n_{1}-1)}\frac{1}{n_{1}(n_{1}-1)}\sum_{l=1}^{n_{1}}\left(V_{l,0,0}-\bar{V}_{\cdot,0,0}\right)^{2}$
$\qquad+\frac{(n_{2}-m_{2})^{2}}{n_{2}(n_{2}-1)}\frac{1}{n_{2}(n_{2}-1)}\sum_{l=n_{1}+1}^{n_{1}+n_{2}}\left(V_{0,l,0}-\bar{V}_{0,\cdot,0}\right)^{2}+\frac{(n_{3}-m_{3})^{2}}{n_{3}(n_{3}-1)}\frac{1}{n_{3}(n_{3}-1)}\left(V_{0,0,l}-\bar{V}_{0,0,\cdot}\right)^{2}\bigg]$
$=n\hat{\sigma}^{2}+o_{p}(1).$
By substituting the above into equation (*) and using Theorem 2.1, we obtain
$S_{n}^{2}=n\sigma^{2}+o_{p}(1).$
This prove the assertion. ∎
The proof of the following Lemma is straight forward and on the same lines as of an2018jackknife, and hence omitted.
Lemma 6.3.
Suppose that conditions (a) and (b) of the Theorem 2.3 hold. Then,
$\max_{1\leq l\leq n}\left|V_{l}-EV_{l}\right|=o_{p}\left(n^{1/2}\right)\;\;\text{ and }\;\;\frac{1}{n}\sum_{l=1}^{n}\left|V_{l}-EV_{l}\right|^{3}=o_{p}\left(n^{1/2}\right).$
Lemma 6.4.
Under the conditions (a) and (b) of the Theorem 2.3, we have
$|\lambda|=O_{p}\left(n^{-1/2}\right)\;\;\text{ and }\;\;\lambda=\frac{\left(U-\theta\right)}{S_{n}^{2}}+o_{p}\left(n^{-1/2}\right).$
Proof.
Noting that, (2.16) together with the fundamental inequality $\mid x\pm$ $y|\geq|x|-|y\mid$ leads to
$0$ $=\left|\frac{1}{n}\sum_{i=1}^{n}\frac{V_{i}-EV_{i}}{1+\lambda\left(V_{i}-EV_{i}\right)}\right|=\frac{1}{n}\left|\sum_{i=1}^{n}\left(V_{i}-EV_{i}\right)-\lambda\sum_{i=1}^{n}\frac{\left(V_{i}-EV_{i}\right)^{2}}{1+\lambda\left(V_{i}-EV_{i}\right)}\right|$
$\geq\frac{1}{n}\left|\lambda\sum_{i=1}^{n}\frac{\left(V_{i}-EV_{i}\right)^{2}}{1+\lambda\left(V_{i}-EV_{i}\right)}\right|-\frac{1}{n}\left|\sum_{i=1}^{n}\left(V_{i}-EV_{i}\right)\right|$
$\geq\frac{|\lambda|S_{n}^{2}}{1+|\lambda|\left(\max_{1\leq l\leq n}\left|V_{l}-EV_{l}\right|\right)}-\left|U_{n}-\theta\right|$
Using the Theorem 2.1, the second term is $O_{p}\left(n_{1}^{-1/2}\right)$ and, from Lemma 6.2, we obtain
$|\lambda|\left(1+|\lambda|\left(\max_{1\leq l\leq n}\left|V_{l}-EV_{l}\right|\right)\right)^{-1}=O_{p}\left(n^{-1/2}\right).$
Therefore, using $\max_{1\leq l\leq n}\left|V_{l}-EV_{l}\right|=o_{p}\left(n^{1/2}\right)$ from Lemma 6.3, we have
$|\lambda|=O_{p}\left(n^{-1/2}\right).$ (6.6)
To prove that second part of lemma, let $\gamma_{l}=\lambda\left(V_{l}-EV_{l}\right),\;l=1,2,\dots,n$. Then, using Lemma 6.3, we get
$\max_{1\leq l\leq n}\left|\gamma_{l}\right|$ $=|\lambda|\max_{1\leq l\leq n}\left|V_{l}-EV_{l}\right|=O_{p}\left(n^{-1/2}\right)o\left(n^{1/2}\right)=o_{p}(1).$ (6.7)
From (2.16), we have
$0$ $=\frac{1}{n}\sum_{i=1}^{n}\frac{V_{i}-EV_{i}}{1+\lambda\left(V_{i}-EV_{i}\right)}$
$=\frac{1}{n}\sum_{i=1}^{n}\left(V_{i}-EV_{i}\right)-\lambda\frac{1}{n}\sum_{i=1}^{n}\frac{\left(V_{i}-EV_{i}\right)^{2}}{1+\lambda\left(V_{i}-EV_{i}\right)}$
$=\frac{1}{n}\sum_{i=1}^{n}\left(V_{i}-EV_{i}\right)-\lambda\frac{1}{n}\sum_{i=1}^{n}\frac{\left(\left(V_{i}-EV_{i}\right)^{2}(1+\gamma_{i})-\lambda\left(V_{i}-EV_{i}\right)^{3}\right)}{1+\gamma_{i}}$
$=U_{n}-\theta-\lambda S_{n}^{2}+\lambda^{2}\frac{1}{n}\sum_{i=1}^{n}\frac{\left(V_{i}-EV_{i}\right)^{3}}{1+\gamma_{i}}.$
Also, using Lemma 6.3 and equations (6.6)-(6.7), last term in the above expression is bounded by
$\frac{\frac{1}{n}\sum_{i=1}^{n}\left|V_{i}-EV_{i}\right|^{3}}{\left|1+\max_{1\leq l\leq n}\left|\gamma_{l}\right|\right|}\lambda^{2}=o_{p}\left(n^{1/2}\right)o_{p}(1)O_{p}\left(n^{-1}\right)=o_{p}\left(n^{-1/2}\right)$
Therefore, we get
$\lambda=\frac{\left(U-\theta\right)}{S_{n}^{2}}+o_{p}\left(n^{-1/2}\right).$
Hence the lemma is proved. ∎
Proof.
of Theorem 2.3: Now, we are ready to prove the theorem. Take $\gamma_{l}=\lambda\left(V_{l}-EV_{l}\right),\;l=1,2,\dots,n,$ and consider the $\log$ likelihood ratio given by
$-2\log R(\theta)$ $=2\sum_{1=1}^{n}\log\left(1+\lambda\left(V_{i}-EV_{i}\right)\right)$
$=2\sum_{1=1}^{n}\log\left(1+\gamma_{i}\right).$
From the Taylor series expansion for $\log\left(1+\gamma_{i}\right)$, we obtain
$2\log R(\theta)=2\sum_{1=1}^{n}\gamma_{i}-\sum_{1=1}^{n}{\gamma_{i}}^{2}+\frac{2}{3}\sum_{1=1}^{n}\gamma_{i}{}^{3}-\frac{1}{2}\sum_{1=1}^{n}{\gamma_{i}}^{4}+\ldots$ (6.8)
Let us first consider the third term of (6.8)
$\frac{2}{3}\sum_{1=1}^{n}\gamma_{i}{}^{3}=\frac{2}{3}\sum_{1=1}^{n}\lambda^{3}\left(V_{i}-EV_{i}\right)^{3}$ $\leq\frac{2}{3}\,n^{\frac{3}{2}}|\lambda|^{3}\frac{1}{n}\sum_{1=1}^{n}\left|V_{i}-EV_{i}\right|^{3}\frac{1}{\sqrt{n}}$
$=n^{\frac{3}{2}}O_{p}(n^{-3/2})o_{p}(n^{\frac{1}{2}})o_{p}(n^{-\frac{1}{2}})$ (By using Lemmas 6.3 and 6.4)
$=o_{p}(1).$
Hence, it is straightforward to verify that all terms with powers of $\gamma_{i}$ equal to or greater than 3 will be $o_{p}(1)$.
Next, we consider the second term of (6.8) as
$\sum_{1=1}^{n}\gamma_{i}^{2}=\sum_{1=1}^{n}\lambda^{2}\left(V_{i}-\theta\right)^{2}=n\lambda^{2}S_{n}^{2}$ $\leq n|\lambda|^{2}(nS_{n_{1},n_{2},n_{3}}^{2}+o_{p}(1))$ (By using Lemma 6.2)
$=n\;o_{p}(n^{-1})(nS_{n_{1},n_{2},n_{3}}^{2}+o_{p}(1))$ (By using Lemma 6.4)
$=o_{p}(1).$
Now, consider the first term of (6.8)
$2\sum_{1=1}^{n}\gamma_{i}=2\sum_{1=1}^{n}\lambda\left(V_{i}-EV_{i}\right)$ $=2\lambda n\frac{1}{n}\sum_{1=1}^{n}\left(V_{i}-EV_{i}\right)$
$=2\lambda n(U_{n}-\theta)$
$=\frac{n(U_{n}-\theta)^{2}}{S_{n}^{2}}+no_{p}(n^{-1/2})(U_{n}-\theta)$ (By using Lemma 6.4)
$=\frac{n(U_{n}-\theta)^{2}}{S_{n}^{2}}+n\,o_{p}(n^{-1/2})O_{p}(n^{-1/2})$ (By using Theorem 2.1)
$=\frac{n(U_{n}-\theta)^{2}}{S_{n}^{2}}+o_{p}(1).$
Therefore,
$-2\log R(\theta)=\frac{n(U_{n}-\theta)^{2}}{S_{n}^{2}}+o_{p}(1)=\frac{(U_{n}-\theta)^{2}}{\sigma^{2}}+o_{p}(1)\;\;\;\;\text{as }\;n_{1}\rightarrow\infty.$ (By using Lemma 6.2)
By using part (b) of Theorem 2.1 and the Slutsky’s theorem, $2\log R(\theta)$ converges in distribution to a $\chi^{2}$ random variable with one degree of freedom. Hence the theorem is proved.